### v3

# Direct and Converse Theorems in Estimating Signals with Sublinear Sparsity

Abstract
This paper addresses the estimation of signals with sublinear sparsity sent over the additive white Gaussian noise channel. This fundamental problem arises in designing denoisers used in message-passing algorithms for sublinear sparsity. From a theoretical perspective, the main results are direct and converse theorems in the sublinear sparsity limit, where the signal sparsity grows sublinearly in the signal dimension as the signal dimension tends to infinity. As a direct theorem, the maximum likelihood estimator is proved to achieve vanishing square error in the sublinear sparsity limit if the noise variance is smaller than a threshold. This threshold is known to be achievable by an existing separable Bayesian estimator. As a converse theorem, all estimators cannot achieve square errors smaller than the signal power under a mild condition if the noise variance is larger than another threshold. In particular, the two thresholds coincide with each other when non-zero signals have constant amplitude. These results imply the asymptotic optimality of the existing separable Bayesian estimator used in approximate message-passing for sublinear sparsity. From a numerical perspective, a non-separable estimator is proposed via a heuristic approximation of the true posterior mean estimator. Numerical simulations show that the ML estimator and the proposed non-separable estimator outperform the separable Bayesian estimator for high signal-to-noise ratio (SNR). In the low SNR regime, on the other hand, the two estimators are inferior to the separable Bayesian estimator while the proposed non-separable estimator slightly outperforms the ML estimator.
publicationid: pubid: 0000–0000/00\$00.00 © xxxx IEEE
I Introduction
I-A Motivation
Estimate an unknown $N$-dimensional and $k$-sparse signal vector $\boldsymbol{x}$ sent over the additive white Gaussian noise (AWGN) channel,
$\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{\omega},\quad\boldsymbol{\omega}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{N}),$ (1)
with noise variance $\sigma_{N/k}^{2}=\sigma^{2}/\log(N/k)$ for some $\sigma^{2}>0$. The signal support $\mathcal{S}_{\boldsymbol{x}}=\{n\in\mathcal{N}=\{1,\ldots,N\}:x_{n}\neq 0\}$ is assumed to be sampled from all possible sets $\mathfrak{S}_{k}^{N}=\{\mathcal{S}\subset\mathcal{N}:|\mathcal{S}|=k\}$ of size $k$ uniformly and randomly. In particular, this paper focuses on signals with sublinear sparsity: The signal sparsity kk—number of non-zero elements in $\boldsymbol{x}$—grows sublinearly in $N$ as $N\to\infty$. The goal is estimation of the sparse signal vector $\boldsymbol{x}$ based on the observation $\boldsymbol{y}$. This simple problem arises in designing denoisers used in message-passing algorithms for sublinear sparsity [1, 2].
The noise power scaling $\sigma_{N/k}^{2}=\sigma^{2}/\log(N/k)$ in (1) is a key point in estimation of the sparse signal vector. Conventional research for sublinear sparsity [3, 4, 5, 6] postulated $N^{-1}\mathbb{E}[\|\boldsymbol{\omega}\|_{2}^{2}]=\Omega(1)$. This normalization implies that error-free estimation is impossible as long as the non-zero signal power is ${\cal O}(1)$. In evaluating the performance of an estimator $\hat{\boldsymbol{x}}\in\mathbb{R}^{N}$, the normalized square error $N^{-1}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{2}^{2}$ was used. For the scaling in this paper, on the other hand, $k^{-1}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|_{2}^{2}$ normalized by sparsity kk—simply called square error—should be considered [1, 2].
We know that the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$ minimizes the mean square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{opt}}\|_{2}^{2}]$. However, the true posterior mean estimator requires high complexity. As a suboptimal low-complexity estimator, the element-wise posterior mean estimator $\mathbb{E}[x_{n}|y_{n}]$—called separable Bayesian estimator—was proposed in [1]. When the essential infimum $u_{\mathrm{min}}$ of non-zero absolute signals $\{|x_{n}|:n\in\mathcal{S}_{\boldsymbol{x}}\}$ is strictly positive, this estimator was proved to achieve asymptotically zero square error if and only if $\sigma^{2}$ is smaller than $u_{\mathrm{min}}^{2}/2$ [1].
When the separable Bayesian estimator is used as a denoiser, both approximate message-passing (AMP) [1] and orthogonal AMP (OAMP) [2] for sublinear sparsity require strictly larger sample complexity than information-theoretically optimal complexity [7, 8, 9, 10, 11, 12, 13]. Early research [7, 8, 9] elucidated the scaling of the optimal complexity for the linear model in compressed sensing while [10, 11] tackled generalized linear models. The prefactor in the optimal scaling was evaluated in [12, 13] for constant non-zero signals. These existing results depend strongly on the assumption of independent and identically distributed (i.i.d.) sensing matrices, so that the linear model in compressed sensing cannot be reduced to the AWGN channel (1). Thus, the existing results are not directly applicable to signal estimation from the observation in the AWGN channel (1).
The suboptimality for sublinear sparsity is in contrast to the optimality of AMP and OAMP for conventional linear sparsity $k/N\to\rho$ for some $\rho\in(0,1]$: For linear sparsity, AMP [14, 15] was proved to be Bayes-optimal [16, 17, 18] via state evolution [19, 20, 21, 22] when sparse signals are compressed with zero-mean i.i.d. sub-Gaussian matrices. Similarly, OAMP [23] or equivalently vector AMP (VAMP) [24] was proved to be Bayes-optimal [25, 26] for right-orthogonally invariant matrices via state evolution [24, 27]. See [28, 29] for a generalization of right-orthogonally invariant matrices. In contrast to these optimality for linear sparsity, both AMP and OAMP are suboptimal for sublinear sparsity [1, 2].
To investigate room for improvement in AMP and OAMP for sublinear sparsity, this paper presents information-theoretic analysis for estimation of the sparse signal vector in the AWGN channel (1). In the analysis, non-zero signals are assumed to be sampled from a discrete set uniformly and randomly. The aims of this simplifying assumption are twofold:
  • In the direct part, i.e. achievability of vanishing square error, Gallager’s bound [30] is available for evaluating the performance of the maximum likelihood (ML) estimator. Gallager’s bound is a general technique that was utilized to prove achievability in sparse linear regression [11]. In contrast to a long proof for the separable Bayesian estimator [1] via a heuristic strategy, the proof of the direct part is simplified for the ML estimator.
  • Sparse signals with discrete non-zero elements can be regarded as index-modulated signals in communication [31, 32, 33, 34]. Thus, OAMP [2] is available for detection of index-modulated signals while AMP [1] requires a strong assumption on the channel matrix. The latter aim is outside the scope of this paper, so that this communication issue is not discussed anymore.
The converse part, i.e. impossibility of vanishing square error is proved via a hypothesis testing [35, 36]. In the hypothesis testing, we test whether the null hypothesis $\boldsymbol{y}\sim q(\boldsymbol{y})$ holds for some probability density function (pdf) $q(\boldsymbol{y})$ different from $p(\boldsymbol{y})$ or the alternative hypothesis $\boldsymbol{y}\sim p(\boldsymbol{y})$. When the true marginal pdf $p(\boldsymbol{y})$ of $\boldsymbol{y}$ in (1) is indistinguishable from $q(\boldsymbol{y})$ independent of the channel model (1), it is intuitively impossible to estimate the signal vector $\boldsymbol{x}$. Inspired by [13], this paper utilizes a general relationship [37] between mutual information and square error to derive a lower bound on the square error, which depends on the Kullback-Leibler (KL) divergence $D(p\|q)$ of $q(\boldsymbol{y})$ from the truth $p(\boldsymbol{y})$. In this paper, a Gaussian pdf $q(\boldsymbol{y})$ is selected. By proving that $D(p\|q)$ converges to zero, this paper proves a converse result.
I-B Contributions
The contributions of this paper are threefold: achievability, converse, and numerical simulations. Part of these contributions were presented in [38].
Before presenting theoretical results of this paper, additional assumptions are summarized. Throughout this paper, non-zero elements of the signal vector are sampled from the set of $M$ discrete non-zero points $\mathcal{U}=\{u_{m}\in\mathbb{R}\setminus\{0\}:m\in\mathcal{M}=\{1,\ldots,M\}\}$ uniformly and randomly. Let $u_{\mathrm{min}}=\min_{m\in\mathcal{M}}|u_{m}|$ and $u_{\mathrm{max}}=\max_{m\in\mathcal{M}}|u_{m}|$. The set of all possible $k$-sparse signal vectors is written as $\mathcal{X}_{k}^{N}(\mathcal{U})=\{\boldsymbol{x}\in\mathbb{R}^{N}:|\mathcal{S}_{\boldsymbol{x}}|=k,x_{n}\in\mathcal{U}~\hbox{for all $n\notin\mathcal{S}_{\boldsymbol{x}}$}\}$.
Theorem 1 (Achievability)
Let $\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ denote the ML estimator of $\boldsymbol{x}$ based on the observation $\boldsymbol{y}$ in (1). If $\sigma^{2}<u_{\mathrm{min}}^{2}/2$ holds, then the square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]$ converges to zero in the sublinear sparsity limit: $N$ tends to infinity with $\log k/\log N\to\gamma$ for some $\gamma\in[0,1)$.
Theorem 1 implies that the ML estimator satisfies the same achievability as the separable Bayesian estimator [1] in the low noise regime $\sigma^{2}<u_{\mathrm{min}}^{2}/2$. See Appendix A-A for a review of the separable Bayesian estimator. Since the ML estimator can be suboptimal in terms of the square error, Theorem 1 makes no claims for high noise regime. This paper presents the following converse result for high noise regime:
Theorem 2 (Converse)
Let $\hat{\boldsymbol{x}}(\boldsymbol{y})\in\mathbb{R}^{N}$ denote any estimator of $\boldsymbol{x}$ given $\boldsymbol{y}$ in (1). If $\sigma^{2}>u_{\mathrm{max}}^{2}/2$ is satisfied, then
$\liminf_{N\to\infty}\frac{1}{k}\left(\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]-\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]\right)\geq 0$ (2)
holds in the sublinear sparsity limit for all
$\gamma<\min\left\{\frac{C_{0}}{C_{0}+\sigma^{2}/(u_{\mathrm{max}}u_{\mathrm{min}})},\;\frac{C_{1}}{C_{1}+3}\right\},$ (3)
with
$C_{0}=\frac{\sigma^{2}}{u_{\mathrm{max}}^{2}}\left(1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}\right),$ (4)
$C_{1}=\frac{\sigma^{2}}{2u_{\mathrm{max}}^{2}}\left(1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}\right)^{2}+\frac{(u_{\mathrm{max}}-u_{\mathrm{min}}/2)^{2}}{2\sigma^{2}}.$ (5)
Theorem 2 claims a stronger result than the strict positivity of the square error: It is impossible to reduce the square error from the prior value $k^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]$. The condition $\sigma^{2}>u_{\mathrm{max}}^{2}/2$ is required to prove this strong converse result. Weak converse results need to be considered to investigate the intermediate regime $\sigma^{2}\in(u_{\mathrm{min}}^{2}/2,u_{\mathrm{max}}^{2}/2]$.
We can obtain a tight result for constant non-zero signals, i.e. $u_{\mathrm{min}}=u_{\mathrm{max}}$. The upper bound in (3) is strictly positive as long as $\sigma^{2}>u_{\mathrm{min}}^{2}/2$ is satisfied. See Fig. 1 for the quantitative evaluation of $\gamma$ satisfying (3). Combining Theorems 1 and 2, this paper concludes the existence of small $\gamma$ such that the square error converges to zero in the sublinear sparsity limit if and only if $\sigma^{2}<u_{\mathrm{min}}^{2}/2$ holds. In other words, the separable Bayesian estimator [1] is asymptotically optimal for small $\gamma$.
Figure 1: Upper bound (3) on $\gamma$ versus $u_{\mathrm{max}}^{2}/\sigma^{2}$. The lower bound (2) in Theorem 2 holds when $(u_{\mathrm{max}}^{2}/\sigma^{2},\gamma)$ is located below each curve.
The last contribution is numerical results. A non-separable estimator is proposed via a heuristic approximation of the posterior mean estimator in the noiseless limit. This estimator is called non-separable approximate Bayesian estimator or more simply non-separable estimator. The proposed non-separable estimator is numerically shown to outperform the ML estimator for all signal-to-noise ratios (SNRs) in terms of the square error. Furthermore, it is superior to the separable Bayesian estimator [1] in the high SNR regime. Unfortunately, the non-separable estimator cannot improve the performance of AMP [1] for sublinear sparsity compared to the separable Bayesian estimator. The last result suggests that AMP for sublinear sparsity has a performance bottleneck in another part rather than the type of denoisers.
I-C Organization
The remainder of this paper is organized as follows: After summarizing the notation used in this paper, Theorem 1 is proved in Section II. The ML estimator is defined in this section. Section III presents the proof of Theorem 2. The two theorems are proved with several technical lemmas proved in Appendices. Section IV presents numerical results for AMP [1]. Section V concludes this paper.
I-D Notation
Throughout this paper, $\boldsymbol{A}^{\mathrm{T}}$ denotes the transpose of a matrix $\boldsymbol{A}$. The notation $\boldsymbol{1}$ represents a vector of which the elements are all one while $1(\cdot)$ denotes the indicator function. The $p$-norm is written as $\|\cdot\|_{p}$. For a vector $\boldsymbol{v}$, its $n$th element is written as $v_{n}$. The notation $\boldsymbol{v}_{\setminus n}$ denotes the vector obtained by eliminating the element $v_{n}$ from $\boldsymbol{v}$. For a set of indices $\mathcal{I}$, the notation $\boldsymbol{v}_{\mathcal{I}}$ represents the vector obtained by extracting the elements $\{v_{n}:n\in\mathcal{I}\}$ from $\boldsymbol{v}$. The $n$th maximum element of a vector $\boldsymbol{v}$ is written as $v_{(n)}$ or $\boldsymbol{v}]_{(n)}$. The notational convention $\sum_{n=i}^{j}\cdots=0$ is used for $i>j$.
The notation $p(X)$ denotes the pdf of an absolutely continuous random variable $X$ while $\mathbb{P}(\cdot)$ represents a probability measure. The Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ is denoted by $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. The function $Q(x)$ represents the complementary cumulative distribution function of the standard Gaussian distribution. See Table I at the end of this paper for the list of the other special notation used throughout this paper.
II Direct Part
II-A Maximum Likelihood Estimation
Before proving Theorem 1, the ML estimator is presented. The ML estimator $\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ of $\boldsymbol{x}$ based on $\boldsymbol{y}$ in (1) is defined as
$\hat{\boldsymbol{x}}_{\mathrm{ML}}=\mathop{\mathrm{argmin}}\limits_{\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})}p(\boldsymbol{y}|\boldsymbol{x}).$ (6)
Since $\boldsymbol{x}$ is uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$, the ML estimator (6) is equivalent to the maximum a posteriori (MAP) estimator
$\hat{\boldsymbol{x}}_{\mathrm{MAP}}=\mathop{\mathrm{argmin}}\limits_{\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})}p(\boldsymbol{x}|\boldsymbol{y}),$ (7)
which minimizes the error probability $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{MAP}}\neq\boldsymbol{x})$. In terms of AMP [1] and OAMP [2] for sublinear sparsity, the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$ should be used to minimize the mean square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{opt}}\|_{2}^{2}]$. Nonetheless, this paper focuses on the ML estimator (6) because the optimality of the ML estimator is partially proved in the converse part.
The ML estimator (6) can be represented via a partition of indices $\mathcal{N}=\{1,\ldots,N\}$: For some $k_{*}\in\{0,\ldots,k\}$ decompose $\mathcal{N}$ into three disjoint sets $\mathcal{N}_{1}\subset\mathcal{N}$, $\mathcal{N}_{2}\subset\mathcal{N}\setminus\mathcal{N}_{1}$, and $\mathcal{N}_{3}=\mathcal{N}\setminus(\mathcal{N}_{1}\cup\mathcal{N}_{2})$ with cardinality $_{*}, kk$$k_{*}$, and NkN-k, respectively. The set $\mathcal{N}_{1}$ consists of indices that correspond to the top $k_{*}$ maximum received signal values while $\mathcal{N}_{2}$ is composed of indices for the top $(k-k_{*})$ minimum received signal values. The set $\mathcal{N}_{3}$ consists of the remaining indices.
Let $\boldsymbol{y}]_{(n)}$ denote the $n$th maximum element of $\boldsymbol{y}$, i.e. $\boldsymbol{y}]_{(1)}\geq\cdots\geq[\boldsymbol{y}]_{(N)}$. Furthermore, we define the permutation $\sigma:\mathcal{N}\to\mathcal{N}$ satisfying $y_{\sigma(n)}=[\boldsymbol{y}]_{(n)}$ for all $n\in\mathcal{N}$. Using these notations, we have the representation $\mathcal{N}_{1}=\{\sigma(n):n\in\{1,\ldots,k_{*}\}\}$, $\mathcal{N}_{2}=\{\sigma(n):n\in\{N-(k-k_{*})+1,\ldots,N\}\}$, and $\mathcal{N}_{3}=\{\sigma(n):n\in\{k_{*}+1,\ldots,N-(k-k_{*})\}\}$.
Proposition 1
Determine $\mathcal{N}_{1}$, $\mathcal{N}_{2}$, and $\mathcal{N}_{3}$ with
$k_{*}=\mathop{\mathrm{argmin}}\limits_{k_{0}\in\{0,\ldots,k\}}\xi_{k}^{N}(k_{0};\boldsymbol{y}),$ (8)
where $\xi_{k}^{N}(k_{0};\boldsymbol{y})$ is given by
$\xi_{k}^{N}(k_{0};\boldsymbol{y})$ $=\sum_{i=1}^{k_{0}}\min_{u\in\mathcal{U}}([\boldsymbol{y}]_{(i)}-u)^{2}+\sum_{i=k_{0}+1}^{N-(k-k_{0})}[\boldsymbol{y}]_{(i)}^{2}$
$+\sum_{i=N-(k-k_{0})+1}^{N}\min_{u\in\mathcal{U}}([\boldsymbol{y}]_{(i)}-u)^{2}.$ (9)
Then, the ML estimator (6) satisfies $\hat{x}_{\mathrm{ML},n}=0$ for all $n\in\mathcal{N}_{3}$. For the remaining elements we have
$\hat{x}_{\mathrm{ML},n}=\mathop{\mathrm{argmin}}\limits_{u\in\mathcal{U}}(y_{n}-u)^{2}$ (10)
for all $n\in\mathcal{N}_{1}\cup\mathcal{N}_{2}$.
Proof:
See Appendix A-B. ∎
The partition $\mathcal{N}=\mathcal{N}_{1}\cup\mathcal{N}_{2}\cup\mathcal{N}_{3}$ can be computed in ${\cal O}(N)$ time. Furthermore, the minimizer (8) does not depend on $\sum_{i=k+1}^{N-k}[\boldsymbol{y}]_{(i)}^{2}$, so that the minimization problem (8) can be solved in ${\cal O}(k)$ time. Thus, the complexity of the ML estimator is linear in $N$.
II-B Proof of Theorem 1
To prove Theorem 1, we focus on the number of incorrectly detected non-zero positions in support recovery and the number of incorrectly detected non-zero signals after support recovery. For the former number, let $w=|\mathcal{S}_{\boldsymbol{x}}\setminus\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}|\in\{0,\ldots,k\}$. The number of correctly detected positions is $|\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}|=k-w$. Furthermore, we have $|\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}\setminus\mathcal{S}_{\boldsymbol{x}}|=w$. When the ML estimator detects (kw)(k-w) positions correctly, the error in support recovery increases the square error by at most $2u_{\mathrm{max}}^{2}w$, with $u_{\mathrm{max}}=\max_{m\in\mathcal{M}}|u_{m}|$.
For the latter number, we write the number of incorrectly detected non-zero elements in correctly detected positions $n\in\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}$ as $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=|\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}|-|\{n\in\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}:x_{n}=\hat{x}_{\mathrm{ML},n}\}|$. When $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}$ holds for $w^{\prime}\in\{0,\ldots,k-w\}$, the error in non-zero signal detection increases the square error by at most $d_{\mathrm{max}}^{2}w^{\prime}$, with $d_{\mathrm{max}}=\max_{m,m^{\prime}\in\mathcal{M}:m\neq m^{\prime}}|u_{m}-u_{m^{\prime}}|$. These two error events imply that the square error is bounded from above by $2u_{\mathrm{max}}^{2}w+d_{\mathrm{max}}^{2}w^{\prime}\leq\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}k$.
We evaluate the square error $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]$, in which the expectation is over $\boldsymbol{x}$ and $\boldsymbol{y}$ in (1). The signal vector $\boldsymbol{x}$ is uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$. Let $\mathcal{E}_{d}=\{\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}\geq\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}d\}$ with $d=\lceil k/\sqrt{\log(N/k)}\rceil$. Since $\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}$ is bounded from above by $\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}k$, we use the representation $\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]=\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}|\mathcal{E}_{d}]\mathbb{P}(\mathcal{E}_{d})+\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}|\mathcal{E}_{d}^{\mathsf{c}}]\mathbb{P}(\mathcal{E}_{d}^{\mathsf{c}})$ to obtain the following upper bound:
$\frac{1}{k}\mathbb{E}\left[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}\right]<\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}\left\{\mathbb{P}(\mathcal{E}_{d})+\frac{d}{k}\right\}.$ (11)
Thus, $\mathbb{P}(\mathcal{E}_{d})\to 0$ implies $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}]\to 0$ in the sublinear sparsity limit.
The main challenge is evaluation of $\mathbb{P}(\mathcal{E}_{d})$. To evaluate this probability, we classify the signal space into types. Let $\boldsymbol{u}_{\boldsymbol{m}}=[u_{m_{1}},\ldots,u_{m_{k}}]^{\mathrm{T}}\in\mathcal{U}^{k}$ denote the vector of non-zero signal elements, with an index vector $\boldsymbol{m}\in\mathcal{M}^{k}$. For a subset $\mathcal{S}\subset\mathfrak{S}_{k}^{N}$ of size $k$, $\boldsymbol{m}\in\mathcal{M}^{|\mathcal{S}|}$, and $w\in\{0,\ldots,|\mathcal{S}|\}$, define the type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})$ of $|\mathcal{S}|$-sparse vectors in $\mathcal{X}_{|\mathcal{S}|}^{N}(\mathcal{U})$ as
$\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})=\left\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{|\mathcal{S}|}^{N}(\mathcal{U}):|\mathcal{S}_{\boldsymbol{x}^{\prime}}\setminus\mathcal{S}|=w,\;\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}^{\prime}}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}\right\}.$ (12)
The type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ in (12) contains all possible $k$-sparse vectors that have non-zero elements $\boldsymbol{u}_{\boldsymbol{m}}$ and $w$ incorrectly detected positions for the signal vector $\boldsymbol{x}$. By definition, we have the following partition of the signal space $\mathcal{X}_{k}^{N}(\mathcal{U})$:
$\mathcal{X}_{k}^{N}(\mathcal{U})=\bigcup_{w=0}^{k}\bigcup_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}).$ (13)
We reduce the evaluation of $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})$ to that of $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$ for $w^{\prime}\in\{0,\ldots,k-w\}$. The integer $w^{\prime}\in\{0,\ldots,k-w\}$ in $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}$ represents the number of incorrectly detected non-zero elements in correctly detected positions $n\in\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}$. Thus, this joint event implies $\|\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{ML}}\|_{2}^{2}\leq 2u_{\mathrm{max}}^{2}w+d_{\mathrm{max}}^{2}w^{\prime}\leq\max\{2u_{\mathrm{max}}^{2},d_{\mathrm{max}}^{2}\}(w+w^{\prime})$. By definition, $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})$ is affected only by the events satisfying $w+w^{\prime}\geq d$ for all $w\in\{0,\ldots,k\}$ and $w^{\prime}\in\{0,\ldots,k-w\}$:
$\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})$
$=$ $\sum_{w+w^{\prime}\geq d}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\cdot\mathbb{P}\left(\left.\mathcal{E}_{d}\right|\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime},\boldsymbol{x}\right)$
$\leq$ $\sum_{w+w^{\prime}\geq d}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x}),$ (14)
where the last inequality follows from the trivial upper bound $\mathbb{P}(\mathcal{E}_{d}|\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime},\boldsymbol{x})\leq 1$.
We use Gallager’s bound [30] to obtain the following result:
Lemma 1
Let $d_{\mathrm{min}}=\min_{m,m^{\prime}\in\mathcal{M}:m\neq m^{\prime}}|u_{m}-u_{m^{\prime}}|>0$ and $u_{\mathrm{min}}=\min_{m\in\mathcal{M}}|u_{m}|$. Then, in the case of $w=0$ we have
$\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{0,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\binom{k}{w^{\prime}}(M-1)^{w^{\prime}}(N/k)^{-\frac{d_{\mathrm{min}}^{2}w^{\prime}}{8\sigma^{2}}}$ (15)
for all $w^{\prime}\in\{0,\ldots,k\}$ and $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$. On the other hand, in the case of $w>0$ we have
$\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq(ek/w)^{w}\binom{k-w}{w^{\prime}}M^{w}(M-1)^{w^{\prime}}(N/k)^{-E_{w,w^{\prime}}(\sigma^{2})}$ (16)
for all $w^{\prime}\in\{0,\ldots,k-w\}$ and $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$, in which $E_{w,w^{\prime}}(\sigma^{2})$ is given by
$E_{w,w^{\prime}}(\sigma^{2})=\left\{\begin{array}[]{cl}E_{w,w^{\prime}}^{(1)}&\hbox{for $\sigma^{2}<u_{\mathrm{min}}^{2}/8$,}\\ E_{w,w^{\prime}}^{(2)}&\hbox{for $\sigma^{2}\in[u_{\mathrm{min}}^{2}/8,u_{\mathrm{min}}^{2}/2),$}\\ E_{w,w^{\prime}}^{(3)}&\hbox{for $\sigma^{2}\in\left[u_{\mathrm{min}}^{2}/2,\frac{u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}}{2w}\right),$}\\ 0&\hbox{for $\sigma^{2}\geq\frac{u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}}{2w}$,}\end{array}\right$. (17)
with
$E_{w,w^{\prime}}^{(1)}(\sigma^{2})=\left(\frac{u_{\mathrm{min}}^{2}}{4\sigma^{2}}-1\right)w+\frac{d_{\mathrm{min}}^{2}}{8\sigma^{2}}w^{\prime},$ (18)
$E_{w,w^{\prime}}^{(2)}(\sigma^{2})=\frac{\{2(u_{\mathrm{min}}^{2}-\sqrt{2\sigma^{2}u_{\mathrm{min}}^{2}})w+d_{\mathrm{min}}^{2}w^{\prime}\}^{2}}{8\sigma^{2}(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})},$ (19)
$E_{w,w^{\prime}}^{(3)}(\sigma^{2})=\frac{\{(u_{\mathrm{min}}^{2}-2\sigma^{2})w+d_{\mathrm{min}}^{2}w^{\prime}\}^{2}}{8\sigma^{2}(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})}.$ (20)
Proof:
See Appendix B. ∎
Consider $w>0$ and $w^{\prime}=0$. The function $E_{w,0}(\sigma^{2})/w$ in (17) corresponds to Gallager’s reliability function while ww—more precisely $w\log(N/k)$—is analogous to the code length. Furthermore, the threshold $u_{\mathrm{min}}^{2}/2$ corresponds to the channel capacity while $u_{\mathrm{min}}^{2}/8$ corresponds to the so-called cutoff rate. When support recovery and signal estimation are considered jointly, in general, $E_{w,w^{\prime}}(\sigma^{2})$ is not linear in $w$ or $w^{\prime}$.
We utilize the following proposition to evaluate the upper bound (14).
Proposition 2
Define $d=\lceil k/\sqrt{\log(N/k)}\rceil$ and let $r_{k,N}$ denote a variable that depends on $k$ and $N$. If $\limsup|r_{k,N}|=0$ is satisfied in the sublinear sparsity limit, then $\sum_{w=d}^{k}r_{k,N}^{w}$ converges to zero in the sublinear sparsity limit.
Proof:
We evaluate the geometric series as
$\sum_{w=d}^{k}r_{k,N}^{w}=r_{k,N}^{d}\frac{1-r_{k,N}^{k-d+1}}{1-r_{k,N}}\to 0,$ (21)
where the last convergence follows from the definition $d=\lceil k/\sqrt{\log(N/k)}\rceil$ and the assumption $\limsup|r_{k,N}|=0$. ∎
We are ready to prove Theorem 1. From the upper bound (11), it is sufficient to prove the convergence $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})\to 0$ in the sublinear sparsity limit.
For $M=1$ we apply Lemma 1 to the upper bound (14). Since (16) for $w^{\prime}\neq 0$ is zero for $M=1$, we use the linearity $E_{w,0}(\sigma^{2})=wE_{1,0}(\sigma^{2})$ to have
$\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})\leq\sum_{w=d}^{k}(ek/d)^{w}(N/k)^{-wE_{1,0}(\sigma^{2})}\to 0,$ (22)
where the last convergence follows from Proposition 2 with $r_{k,N}=(ek/d)(N/k)^{-E_{1,0}(\sigma^{2})}$, which satisfies
$\log r_{k,N}<-E_{1,0}(\sigma^{2})\log(N/k)+\frac{1}{2}\log\log(N/k)+1\to-\infty,$ (23)
because of $d=\lceil k/\sqrt{\log(N/k)}\rceil$ and the strict positivity of $E_{1,0}(\sigma^{2})$ under the assumption $\sigma^{2}<u_{\mathrm{min}}^{2}/2$. See Appendix C for the proof of $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})\to 0$ in the case of $M>1$.
III Converse Part
III-A Proof Overview
The proof of Theorem 2 consists of two steps: A first step is a hypothesis testing with respect to the distribution of $\boldsymbol{y}$ in (1). As a null hypothesis, the distribution is selected to that of $\boldsymbol{y}$ in (1) with the signal vector $\boldsymbol{x}$ replaced by capacity-achieving Gaussian signaling. The alternative hypothesis is the true distribution. The two distributions are proved to be asymptotically indistinguishable from each other in terms of the KL divergence.
In the second step, we connect the KL divergence to the square error. Inspired by [13], we utilize a general relationship [37] between mutual information and square error to derive a lower bound on the square error. The lower bound depends on the KL divergence and is useful only when the distribution in the null hypothesis is the capacity-achieving output distribution and indistinguishable from the true distribution. Since these conditions are justified in the first step, we use the lower bound to prove Theorem 2.
III-B Hypothesis Testing
A hypothesis testing is considered. For some pdf $q(\boldsymbol{y})$ on $\mathbb{R}^{N}$ different from the true marginal pdf $p(\boldsymbol{y})$ of $\boldsymbol{y}$, consider the null hypothesis $\boldsymbol{y}\sim q(\boldsymbol{y})$ for the received vector $\boldsymbol{y}$ in (1). The alternative hypothesis is the true marginal pdf $\boldsymbol{y}\sim p(\boldsymbol{y})$. Intuitively, it is impossible to estimate the signal vector $\boldsymbol{x}$ from the received vector $\boldsymbol{y}$ if the true pdf $p(\boldsymbol{y})$ is asymptotically indistinguishable from $q(\boldsymbol{y})$. The difference between $p(\boldsymbol{y})$ and $q(\boldsymbol{y})$ is measured with the KL divergence.
Definition 1
The KL divergence of $q(\boldsymbol{y})$ from the true pdf $p(\boldsymbol{y})$ is defined as
$D(p(\boldsymbol{y})\|q(\boldsymbol{y}))=\int p(\boldsymbol{y})\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}d\boldsymbol{y}.$ (24)
The KL divergence is simply written as $D(p\|q)$.
As the null hypothesis, we select the zero-mean Gaussian pdf $q(\boldsymbol{y})$ with covariance $N^{-1}\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$, in which $\mathbb{E}_{p}$ denotes the expectation with respect to the true pdf $p(\boldsymbol{y})$. We know that this pdf $q(\boldsymbol{y})$ is the capacity-achieving output pdf for the AWGN channel (1[39]. This property is important to prove Theorem 2 in the second step. Since the expectation $\mathbb{E}_{q}[\|\boldsymbol{y}\|_{2}^{2}]$ with respect to $q(\boldsymbol{y})$ is equal to the true expectation $\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]$, we are not allowed to distinguish $p(\boldsymbol{y})$ from $q(\boldsymbol{y})$ via the power difference.
Lemma 2
Suppose $\sigma^{2}>u_{\mathrm{max}}^{2}/2$. For the zero-mean Gaussian pdf $q(\boldsymbol{y})$ with covariance $N^{-1}\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$, the KL divergence $D(p\|q)$ converges to zero in the sublinear sparsity limit for
$\gamma<\min\left\{\frac{C_{0}}{C_{0}+\sigma^{2}/(u_{\mathrm{max}}u_{\mathrm{min}})},\frac{C_{1,1}+C_{1,2}}{C_{1,1}+C_{1,2}+3}\right\},$ (25)
with
$C_{0}=\frac{\sigma^{2}}{u_{\mathrm{max}}^{2}}\left(1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}\right),\quad C_{1,1}=\frac{\sigma^{2}}{2u_{\mathrm{max}}^{2}}\left(1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}\right)^{2},$ (26)
$C_{1,2}=\frac{(u_{\mathrm{max}}-u_{\mathrm{min}}/2)^{2}}{2\sigma^{2}}.$ (27)
Proof:
See Appendix D. ∎
The pdf $q(\boldsymbol{y})$ in Lemma 2 is the marginal pdf of $\boldsymbol{y}$ in the AWGN channel (1) for capacity-achieving Gaussian signaling $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{0},N^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}])$. Lemma 2 implies that, in the high noise regime $\sigma^{2}>u_{\mathrm{max}}^{2}/2$, it is impossible to judge from the observation $\boldsymbol{y}$ whether sparse signaling has been used or Gaussian signaling. In other words, the prior information on signal sparsity is useless in the high noise regime.
The condition $\sigma^{2}>u_{\mathrm{max}}^{2}/2$ cannot be weakened. Extreme value theory [40, p. 302] implies that the expected maximum of $N$ independent zero-mean Gaussian noise samples with variance $\sigma_{N/k}^{2}$ is $\sqrt{2\sigma_{N/k}^{2}\log N}\{1+o(1)\}=\sqrt{2\sigma^{2}(1-\gamma)^{-1}}$ $\cdot\{1+o(1)\}$. For $\sigma^{2}<u_{\mathrm{max}}^{2}/2$, thus, there is small $\gamma>0$ such that the maximum non-zero signal is strictly larger than all Gaussian noise samples. In this case, it is possible to detect the positions of the maximum non-zero signals from the observation $\boldsymbol{y}$.
The condition (25) for $\gamma$ may be weakened. The former condition in $\min\{\cdots\}$ is required to control typical Gaussian noise samples while the latter condition is associated with non-typical noise samples. These two conditions are essential in the proof strategy of Lemma 2. The condition (25) for $\gamma$ may be weakened by using another sophisticated proof strategy.
III-C Square Error
We prove a lemma to connect the KL divergence $D(p\|q)$ with the square error. For notational convenience, we write the conditional pdf $p(\boldsymbol{y}|\boldsymbol{x})$ representing the AWGN channel (1) as $p_{\boldsymbol{y}|\boldsymbol{x}}$. This notation is useful in utilizing the permutation invariance of pdfs. We know that the capacity of the AWGN channel (1)—the supremum of the mutual information $I(\boldsymbol{x};\boldsymbol{y})$ over the distribution of $\boldsymbol{x}$ under the average power constraint $N^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]$—is given by
$C_{\mathrm{AWGN}}=\frac{1}{2}\log\left(1+\frac{\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}{N\sigma_{N/k}^{2}}\right),$ (28)
which is achieved by $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{0},N^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]\boldsymbol{I}_{N})$. In this case, the capacity-achieving output pdf is equal to the zero-mean Gaussian pdf with covariance $N^{-1}\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$, i.e. $q(\boldsymbol{y})$ in Lemma 2.
For this Gaussian pdf $q(\boldsymbol{y})$, define
$J(q(\boldsymbol{y}))$ $=C_{\mathrm{AWGN}}-\mathbb{E}\left[\log\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})}{q(\boldsymbol{y})}\right]+D(p(\boldsymbol{y})\|q(\boldsymbol{y}))$
$=D(p(\boldsymbol{y})\|q(\boldsymbol{y})),$ (29)
where the second equality holds because $q(\boldsymbol{y})$ is equal to the capacity-achieving output pdf for the AWGN channel (1). The following lemma is proved for the capacity-achieving output pdf $q(\boldsymbol{y})$ and the AWGN channel (1), while it is possible to generalize $q(\boldsymbol{y})$ and the conditional pdf $p_{\boldsymbol{y}|\boldsymbol{x}}$ to any pdf and any permutation-invariant conditional pdf, i.e. $p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{P}\boldsymbol{y}|\boldsymbol{P}\boldsymbol{x})=p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})$ for any permutation matrix $\boldsymbol{P}$.
Lemma 3
Let $\hat{\boldsymbol{x}}(\boldsymbol{y})\in\mathbb{R}^{N}$ denote any estimator of $\boldsymbol{x}$ given $\boldsymbol{y}$ in the AWGN channel (1). Then, for the zero-mean Gaussian pdf $q(\boldsymbol{y})$ with covariance $N^{-1}\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$ we have
$\frac{1}{k}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq\frac{1}{\alpha}\left\{C_{N}(\alpha)e^{-2D(p(\boldsymbol{y})\|q(\boldsymbol{y}))}-1\right\}$ (30)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$, with
$C_{N}(\alpha)=\frac{(1-u_{\mathrm{max}}^{4}\alpha^{2})e^{\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}}{\{1+f(\alpha)[(1-k/N)^{-k}-1]\}^{2}},$ (31)
$f(\alpha)=\left(\frac{1-u_{\mathrm{max}}^{4}\alpha^{2}}{1-2u_{\mathrm{max}}^{2}\alpha}\right)^{1/2}-1.$ (32)
Proof:
See Appendix G. ∎
Lemma 3 only provides a trivial bound when $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))$ tends to infinity in the sublinear sparsity limit. In this case, the lower bound (30) reduces to $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq-\alpha^{-1}$ in the sublinear sparsity limit. When $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))\to 0$ holds in the sublinear sparsity limit, on the other hand, we have
$\frac{1}{k}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq\frac{1}{\alpha}\left\{(1-u_{\mathrm{max}}^{4}\alpha^{2})e^{\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}-1\right\}+o(1)$ (33)
in the sublinear sparsity limit under the additional assumption $\gamma<1/2$. Taking the limit $\alpha\to 0$ results in the non-trivial lower bound $k^{-1}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\geq k^{-1}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]+o(1)$.
Proof:
Let $q(\boldsymbol{y})$ denote the zero-mean Gaussian pdf with covariance $N^{-1}\mathbb{E}[\|\boldsymbol{y}\|_{2}^{2}]\boldsymbol{I}_{N}$. Lemma 2 implies $D(p(\boldsymbol{y})\|q(\boldsymbol{y}))\to 0$ in the sublinear sparsity limit under the assumptions in Theorem 2. Under the additional assumption $\gamma<1/2$, Lemma 3 implies Theorem 2. Thus, it is sufficient to confirm that the assumption (3) in Theorem 2 implies $\gamma<1/2$.
Under the change of variables $u=u_{\mathrm{max}}/\sigma$ and $v=u_{\mathrm{min}}/\sigma$, we write the first argument of the minimum operator in the assumption (3) as
$\gamma_{1}(u,v)=\frac{C_{0}}{C_{0}+(uv)^{-1}},\quad C_{0}=\frac{1}{u^{2}}\left(1-\frac{u^{2}}{2}\right).$ (34)
We prove $\gamma_{1}(u,v)<1/2$ for all $u\in(0,\sqrt{2})$ and $v\in(0,u]$. From this inequality, we find that the assumption (3) implies $\gamma<1/2$. Since $\gamma_{1}(u,v)$ is monotonically increasing with respect to $v\in(0,u]$ for fixed $u\in(0,\sqrt{2})$, we obtain $\gamma_{1}(u,v)\leq\gamma_{1}(u,u)\equiv\gamma_{1}(u)$, given by
$\gamma_{1}(u)=\frac{C_{0}u^{2}}{C_{0}u^{2}+1}=\frac{2-u^{2}}{4-u^{2}}<\frac{1}{2}$ (35)
for all $u^{2}\in(0,2)$. Thus, $\gamma_{1}(u,v)<1/2$ holds for all $u\in(0,\sqrt{2})$ and $v\in(0,u]$. ∎
IV Numerical Results
IV-A Non-Separable Approximate Bayesian Estimator
As shown shortly, the ML estimator (6) in Proposition 1 outperforms the separable Bayesian estimator [1] in the low noise regime. However, it is inferior to the separable Bayesian estimator in the high noise regime. To improve the performance of the ML estimator in the high noise regime, we propose the following non-separable approximate Bayesian estimator of $x_{n}$—simply called non-separable estimator:
$\hat{x}_{n}=\frac{\sum_{m\in\mathcal{M}}u_{m}e^{u_{m}y_{n}/\sigma_{N/k}^{2}-\ell_{m}(\boldsymbol{y}_{\setminus n})/(2\sigma_{N/k}^{2})}}{1+\sum_{m\in\mathcal{M}}e^{u_{m}y_{n}/\sigma_{N/k}^{2}-\ell_{m}(\boldsymbol{y}_{\setminus n})/(2\sigma_{N/k}^{2})}},$ (36)
where $\ell_{m}(\boldsymbol{y}_{\setminus n})$ is given by
$\ell_{m}(\boldsymbol{y}_{\setminus n})$ $=u_{m}^{2}+\min_{k_{0}\in\{0,\ldots,k-1\}}\xi_{k-1}^{N-1}(k_{0};\boldsymbol{y}_{\setminus n})$
$-\min_{k_{0}\in\{0,\ldots,k\}}\xi_{k}^{N-1}(k_{0};\boldsymbol{y}_{\setminus n}),$ (37)
with $\xi_{k}^{N}(k_{0};\boldsymbol{y})$ given in (9). See Appendix A-C for a heuristic derivation of (36).
Figure 2: Square error versus SNR $1/\sigma^{2}$ in dB for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, and the AWGN channel (1). $10^{5}$ independent trials were simulated for the non-separable estimator (36), separable Bayesian estimator [1], and ML estimator (6). The vertical dashed line shows the threshold $1/\sigma^{2}=2$ in Theorems 1 and 2.
The constant non-zero signals $\mathcal{U}=\{1\}$ were considered in numerical simulations to compare the three estimators for the AWGN channel (1). Figure 2 shows the square error normalized by $k$ versus SNR $1/\sigma^{2}$. The ML estimator outperforms the separable Bayesian estimator [1] in the high SNR regime while it is inferior to the separable Bayesian estimator in the low SNR regime. The non-separable estimator (36) outperforms the ML estimator (6) for all SNRs. However, it is inferior to the separable Bayesian estimator in the low SNR regime. The high and low SNR regimes are separated at approximately 6 dB. The suboptimality of the non-separable estimator (36) is due to the max-log approximation in the high SNR regime.
As shown in [1, Fig. 1], the sublinear sparsity limit is too far for the threshold $1/\sigma^{2}=2$ to predict a sharp result for finite $N$. The threshold predicts a rough position around which the square error decreases rapidly.
IV-B Compressed Sensing
The separable Bayesian estimator and the non-separable estimator (36) are applied to AMP [1] for compressed sensing with sublinear sparsity. The ML estimator is not considered, because of its non-differentiability. In compressed sensing, the $M$-dimensional measurement vector $\boldsymbol{y}\in\mathbb{R}^{M}$ is given by
$\boldsymbol{y}=k^{-1/2}\boldsymbol{A}\boldsymbol{x}+\boldsymbol{w},\quad\boldsymbol{w}\in\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I}_{N}),$ (38)
where the sensing matrix $\boldsymbol{A}\in\mathbb{R}^{M\times N}$ has independent standard Gaussian elements. In the sublinear sparsity limit, $M/\{k\log(N/k)\}\to\delta>0$ should hold while $\log k/\log N\to\gamma\in[0,1)$ is satisfied as $k$ and $N$ tend to infinity.
AMP using the separable Bayesian estimator was proved to achieve vanishing square error in the sublinear sparsity limit [1, Proposition 1] if $\delta$ is larger than the threshold $\delta_{\mathrm{AMP}}=2(1+\sigma^{2})$. This threshold $\delta_{\mathrm{AMP}}$ is strictly larger than the optimal threshold [13] $\delta_{\mathrm{opt}}=2/\log(1+\sigma^{-2})$.
To improve the convergence property of AMP for finite-sized systems, AMP with damping [1] was simulated. The common damping factor for all iterations was optimized via exhaustive search.
Figure 3: Square errors of AMP using the separable Bayesian estimator and the non-separable estimator versus $\delta=M/\{k\log(N/k)\}$ for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, $1/\sigma^{2}=40$ dB, 3030 iterations, and the measurement model (38). According to the numerical simulations shown in Fig. 2, in non-separable AMP, the non-separable estimator (36) was used when the input SNR to the denoiser is larger than 6 dB. Otherwise, the separable Bayesian estimator was used. $10^{5}$ independent trials were simulated. The optimal threshold [13] is approximately $\delta_{\mathrm{opt}}\approx 0.22$ while the threshold for AMP [1] is $\delta_{\mathrm{AMP}}\approx 2.0$.
Figure 3 shows the square errors for AMP using the separable Bayesian estimator and the non-separable estimator (36). The square errors reduce rapidly as $\delta$ increases toward the threshold $\delta_{\mathrm{AMP}}\approx 2$ for AMP [1]. In contrast to the superiority of the non-separable estimator in the high SNR regime, the two AMP algorithms are comparable to each other for all $\delta$.
To understand this unexpected result, we focus on the input-output relationship for denoising shown in Fig. 4. The iteration trajectories for AMP run from the upper right toward the lower left. If perfect Gaussianity held for estimation errors before denoising in AMP, the two trajectories for AMP would overlap the corresponding input-output relationship for the separable Bayesian estimator and the non-separable estimator. However, AMP outputs much worse square errors than the corresponding estimators for small input square errors while it overlaps the separable Bayesian estimator for large input square errors. Furthermore, AMP using the non-separable estimator is trapped in worse input square errors than the separable counterpart on the final stage of convergence. As a result, the two AMP algorithms achieve almost the same performance.
These observations suggest that the performance bottleneck is in non-ideal properties for estimation errors before denoising, rather than the type of denoisers. Because of finite size effect, the estimation errors cannot be regarded as zero-mean i.i.d. Gaussian random variables. This non-Gaussianity is a dominant factor for the degradation of the convergence property in finite-sized systems.
Figure 4: Square errors after denoising versus the square error normalized by NN before denoising for $N=2^{16}$, $k=16$, constant non-zero signals $\mathcal{U}=\{1\}$, and $1/\sigma^{2}=40$ dB. The measurement model (38) was postulated for the separable/non-separable AMP while the AWGN channel (1) was assumed for the separable Bayesian/non-separable estimators. In non-separable AMP, the non-separable estimator (36) was used when the input SNR to the denoiser is larger than 6 dB. Otherwise, the separable Bayesian estimator was used. $10^{5}$ independent trials were simulated.
V Conclusions
This paper has proved a sharp result for the estimation of signals with sublinear sparsity sent over the AWGN channel: For non-zero signals with unit amplitude, the ML estimator achieves vanishing square error in the sublinear sparsity limit if $\sigma^{2}$ in the noise variance is smaller than $1/2$. Conversely, all estimators result in square errors larger than or equal to $1$ in the sublinear sparsity limit with small $\gamma$ if $\sigma^{2}$ is larger than $1/2$. These theoretical results prove the optimality of the separable Bayesian estimator [1] in the sublinear sparsity limit for small $\gamma$.
Numerical results for AMP suggest that AMP for sublinear sparsity has a performance bottleneck in another part—such as finite size effect—rather than the type of denoisers. To improve the convergence property of AMP, we need to deal with non-Gaussianity for estimation errors before denoising.
Appendix A Estimators
A-A Separable Bayesian Estimator
[…]
Let $U\in\mathbb{R}\setminus\{0\}$ denote a random variable representing i.i.d. non-zero signals. As derived in [1, Section IV-A], the element-wise posterior mean estimator $\mathbb{E}[x_{n}|y_{n}]$ is given by
$\mathbb{E}[x_{n}|y_{n}]=\frac{k\mathbb{E}_{U}[e^{-(y_{n}-U)^{2}/(2\sigma_{N/k}^{2})}]\mathbb{E}[U|U+\omega_{n}=y_{n}]}{k\mathbb{E}_{U}[e^{-(y_{n}-U)^{2}/(2\sigma_{N/k}^{2})}]+(N-k)e^{-y_{n}^{2}/(2\sigma_{N/k}^{2})}},$ (39)
with $\omega_{n}\sim\mathcal{N}(0,\sigma_{N/k}^{2})$ independent of $U$. For the constant non-zero signals $\mathbb{P}(U=1)=1$, the element-wise posterior mean estimator reduces to
$\mathbb{E}[x_{n}|y_{n}]=\frac{ke^{-(y_{n}-1)^{2}/(2\sigma_{N/k}^{2})}}{ke^{-(y_{n}-1)^{2}/(2\sigma_{N/k}^{2})}+(N-k)e^{-y_{n}^{2}/(2\sigma_{N/k}^{2})}}.$ (40)
To compare the main results in this paper with an existing result [1, Lemma 1] for the separable Bayesian estimator, we focus on discrete non-zero signals $\mathcal{U}=\{u_{m}\in\mathbb{R}\setminus\{0\}:m\in\mathcal{M}\}$. We do not necessarily assume the uniform distribution on $\mathcal{U}$. The separable Bayesian estimator (39) was proved to achieve the following square error in the sublinear sparsity limit:
Theorem 3 ([1])
For all $\sigma^{2}>0$, we have
$\frac{1}{k}\sum_{n\in\mathcal{N}}\mathbb{E}\left[\left(x_{n}-\mathbb{E}[x_{n}|y_{n}]\right)^{2}\right]\to\mathbb{E}\left[U^{2}1(U^{2}<2\sigma^{2})\right]$ (41)
in the sublinear sparsity limit.
Proof:
Theorem 3 is equivalent to [1, Lemma 1]. More precisely, [1, Lemma 1] claimed the convergence in probability to treat random small perturbation. When the random small perturbation is eliminated, however, the convergence in mean is what was actually proved in [1].
To use [1, Lemma 1], we need to confirm Assumption 7 in [1, p. 4608]. For the discrete random variable $U\in\mathcal{U}$, Assumption 7 is trivial with the exception of the uniform Lipschitz-continuity for $\mathbb{E}[U|U+\omega_{n}=y_{n}]$ with respect to $\sigma_{N/k}^{2}>0$. In fact, this uniform Lipschitz-continuity is broken for discrete $U\in\mathcal{U}$ as $\sigma_{N/k}^{2}\to 0$. Nonetheless, we can still use [1, Lemma 1]: The uniform Lipschitz-continuity assumption was used in [1] only to justify that there are some constants $L>0$ and $C>0$ such that $|\mathbb{E}[U|U+\omega_{n}=y_{n}]|\leq L|y_{n}|+C$ is satisfied for all $y_{n}\in\mathbb{R}$. Since $\mathbb{E}[U|U+\omega_{n}=y_{n}]$ is uniformly bounded for discrete $U\in\mathcal{U}$, the inequality to be justified is trivial for discrete $U\in\mathcal{U}$. ∎
Theorem 3 implies that the separable Bayesian estimator achieves vanishing square error for all $\sigma^{2}<u_{\mathrm{min}}^{2}/2$, as proved in Theorem 1 for the ML estimator. For all $\sigma^{2}>u_{\mathrm{max}}^{2}/2$, on the other hand, the square error is equal to $\mathbb{E}[U^{2}]$. The latter result is consistent to Theorem 2. In the intermediate region $2\sigma^{2}\in(u_{\mathrm{min}}^{2},u_{\mathrm{max}}^{2})$, Theorem 3 indicates that the square error is in the open interval $(0,\mathbb{E}[U^{2}])$.
A-B Proof of Proposition 1
We represent the $k$-sparse vector $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ with the support set $\mathcal{S}_{\boldsymbol{x}}\in\mathcal{X}_{k}^{N}(\{1\})$ and the non-zero values $\boldsymbol{u}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}\in\mathcal{U}^{k}$. Using the definition of $\boldsymbol{y}$ in (1), we write the ML estimator (6) as
$(\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}},\hat{\boldsymbol{u}})=\mathop{\mathrm{argmin}}\limits_{(\mathcal{S},\boldsymbol{u})\in\mathcal{X}_{k}^{N}(\{1\})\times\mathcal{U}^{k}}\left\{\|\boldsymbol{y}_{\mathcal{S}}-\boldsymbol{u}\|_{2}^{2}+\sum_{n\notin\mathcal{S}}y_{n}^{2}\right\},$ (42)
where $\hat{\boldsymbol{u}}$ corresponds to non-zero elements in the ML estimator.
Let $\boldsymbol{u}_{+}$ and $\boldsymbol{u}_{-}$ denote the vectors obtained by extracting the positive and negative elements from $\boldsymbol{u}$, respectively. For the dimension $k_{0}\in\{0,\ldots,k\}$ of $\boldsymbol{u}_{+}$, we decompose $\mathcal{U}^{k}$ into $k+1$ disjoint subsets $\mathcal{U}^{k}(k_{0})=\{\boldsymbol{u}\in\mathcal{U}^{k}:\boldsymbol{u}_{+}\in\mathcal{U}_{+}^{k_{0}}\}$ with $\mathcal{U}_{+}=\{u_{m}\in\mathcal{U}:u_{m}>0\}$ and $\mathcal{U}_{-}=\{u_{m}\in\mathcal{U}:u_{m}<0\}$. We represent the objective function in the minimization problem (42) as
$\|\boldsymbol{y}_{\mathcal{S}}-\boldsymbol{u}\|_{2}^{2}+\sum_{n\notin\mathcal{S}}y_{n}^{2}$
$=\|\boldsymbol{u}\|_{2}^{2}-2\boldsymbol{u}_{+}^{\mathrm{T}}\boldsymbol{y}_{\mathcal{S}_{+}}-2\boldsymbol{u}_{-}^{\mathrm{T}}\boldsymbol{y}_{\mathcal{S}_{-}}+\|\boldsymbol{y}\|_{2}^{2},$ (43)
where $\mathcal{S}_{\pm}$ consists of the indices of $\boldsymbol{y}_{\mathcal{S}}$ that correspond to $\boldsymbol{u}_{\pm}$. Minimizing this expression over the support sets $\mathcal{S}_{+}$ and $\mathcal{S}_{-}$, we find that the minimizers $\mathcal{S}_{+}^{*}$ and $\mathcal{S}_{-}^{*}$ are the indices that correspond to the top $k_{0}$ maximum and top $(k-k_{0})$ minimum elements of $\boldsymbol{y}$, respectively. Thus, $\hat{x}_{\mathrm{ML},n}=0$ holds for all $n\in\mathcal{N}\setminus(\mathcal{N}_{1}\cup\mathcal{N}_{2})$, with $\mathcal{N}_{1}=\mathcal{S}_{+}^{*}$ and $\mathcal{N}_{2}=\mathcal{S}_{-}^{*}$.
We focus on the minimization over the remaining vector $\boldsymbol{u}$. Let $y_{(n)}$ denote the $n$th maximum of $\{y_{n}:n\in\mathcal{N}\}$. The objective function in (42) minimized over $\mathcal{S}$ is represented as
$\sum_{i=1}^{k_{0}}(y_{(i)}-u_{i})_{2}^{2}+\sum_{i=k_{0}+1}^{N-(k-k_{0})}y_{(i)}^{2}+\sum_{i=N-(k-k_{0})+1}^{N}(y_{(i)}-u_{i})^{2},$ (44)
where $\{u_{i}\in\mathcal{U}_{+}:i\in\{1,\ldots,k_{0}\}\}$ and $\{u_{i}\in\mathcal{U}_{-}:i\in\{N-(k-k_{0})+1,\ldots,N\}\}$ denote the non-zero elements associated with $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, respectively. Minimizing this expression over $\boldsymbol{u}\in\mathcal{U}^{k}(k_{0})$, we find
$\hat{x}_{\mathrm{ML},n}=\mathop{\mathrm{argmin}}\limits_{u\in\mathcal{U}_{+}}(y_{n}-u)^{2}$ (45)
for all $n\in\mathcal{N}_{1}$ and
$\hat{x}_{\mathrm{ML},n}=\mathop{\mathrm{argmin}}\limits_{u\in\mathcal{U}_{-}}(y_{n}-u)^{2}$ (46)
for all $n\in\mathcal{N}_{2}$. Furthermore, we minimize the objective function over $k_{0}\in\{0,\ldots,k\}$ to obtain
$\|\boldsymbol{y}-\boldsymbol{x}\|_{2}^{2}\geq\min_{k_{0}\in\{0,\ldots,k\}}\left\{\sum_{i=1}^{k_{0}}\min_{u\in\mathcal{U}_{+}}([\boldsymbol{y}]_{(i)}-u)^{2}\right$.
$\left.+\sum_{i=k_{0}+1}^{N-(k-k_{0})}[\boldsymbol{y}]_{(i)}^{2}+\sum_{i=N-(k-k_{0})+1}^{N}\min_{u\in\mathcal{U}_{-}}([\boldsymbol{y}]_{(i)}-u)^{2}\right\}.$ (47)
We are ready to prove Proposition 1. By relaxing the feasible sets $\mathcal{U}_{\pm}$ to $\mathcal{U}$, we find that $\xi_{k}^{N}(k_{*},\boldsymbol{y})$ in (9) with $k_{0}=k_{*}$ in (8) is a lower bound on the objective function (47). Since $\xi_{k}^{N}(k_{*},\boldsymbol{y})$ is achievable, we arrive at Proposition 1.
A-C Non-Separable Approximate Bayesian Estimator
We present a heuristic derivation for the non-separable approximate Bayesian estimator (36). We start with the optimum posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$, given by
$\hat{x}_{\mathrm{opt},n}=\frac{\mathbb{E}_{\boldsymbol{x}}[x_{n}p(\boldsymbol{y}|\boldsymbol{x})]}{\mathbb{E}_{\boldsymbol{x}}[p(\boldsymbol{y}|\boldsymbol{x})]},$ (48)
where the expectation is over $\boldsymbol{x}$ uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$. Using the definition of the AWGN channel (1) yields
$\hat{x}_{\mathrm{opt},n}=\frac{\mathbb{E}_{\boldsymbol{x}}[x_{n}e^{-\|\boldsymbol{y}-\boldsymbol{x}\|_{2}^{2}/(2\sigma_{N/k}^{2})}]}{\mathbb{E}_{\boldsymbol{x}}[e^{-\|\boldsymbol{y}-\boldsymbol{x}\|_{2}^{2}/(2\sigma_{N/k}^{2})}]}$
$=\frac{\sum_{m\in\mathcal{M}}u_{m}e^{-\frac{(y_{n}-u_{m})^{2}}{2\sigma_{N/k}^{2}}}}{\sum_{m\in\mathcal{M}}e^{-\frac{(y_{n}-u_{m})^{2}}{2\sigma_{N/k}^{2}}}+e^{-\frac{y_{n}^{2}}{2\sigma_{N/k}^{2}}}e^{L_{\mathrm{opt}}(\boldsymbol{y}_{\setminus n})}},$ (49)
with
$L_{\mathrm{opt}}(\boldsymbol{y}_{\setminus n})=\log\left\{\sum_{\boldsymbol{x}_{\setminus n}\in\mathcal{X}_{k}^{N-1}(\mathcal{U})}e^{-\|\boldsymbol{y}_{\setminus n}-\boldsymbol{x}_{\setminus n}\|_{2}^{2}/(2\sigma_{N/k}^{2})}\right\}$
$-\log\left(\sum_{\boldsymbol{x}_{\setminus n}\in\mathcal{X}_{k-1}^{N-1}(\mathcal{U})}e^{-\|\boldsymbol{y}_{\setminus n}-\boldsymbol{x}_{\setminus n}\|_{2}^{2}/(2\sigma_{N/k}^{2})}\right).$ (50)
To reduce the complexity of $L_{\mathrm{opt}}(\boldsymbol{y}_{\setminus n})$, we use the max-log approximation in the low noise limit $\sigma^{2}\to 0$ to obtain $L_{\mathrm{opt}}(\boldsymbol{y}_{\setminus n})\approx L(\boldsymbol{y}_{\setminus n})$, with
$2\sigma_{N/k}^{2}L(\boldsymbol{y}_{\setminus n})$ $=\min_{\boldsymbol{x}_{\setminus n}\in\mathcal{X}_{k-1}^{N-1}(\mathcal{U})}\|\boldsymbol{y}_{\setminus n}-\boldsymbol{x}_{\setminus n}\|_{2}^{2}$
$-\min_{\boldsymbol{x}_{\setminus n}\in\mathcal{X}_{k}^{N-1}(\mathcal{U})}\|\boldsymbol{y}_{\setminus n}-\boldsymbol{x}_{\setminus n}\|_{2}^{2}.$ (51)
Using Proposition 1 yields
$2\sigma_{N/k}^{2}L(\boldsymbol{y}_{\setminus n})$ $=\min_{k_{0}\in\{0,\ldots,k-1\}}\xi_{k-1}^{N-1}(k_{0};\boldsymbol{y}_{\setminus n})$
$-\min_{k_{0}\in\{0,\ldots,k\}}\xi_{k}^{N-1}(k_{0};\boldsymbol{y}_{\setminus n}),$ (52)
with $\xi_{k}^{N}(k_{0};\boldsymbol{y})$ in (9). Thus, (36) is obtained.
Appendix B Proof of Lemma 1
B-A Overview
The proof of Lemma 1 consists of three steps. A first step is the decomposition of the type $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ in (12) into two types: A first type is associated with the true signal support $\mathcal{S}_{\boldsymbol{x}}$. The other type corresponds to the remaining indices $\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}$. For the former type, the simple union bound is used.
A second step is Gallager’s bound for the latter type. In this step, we follow the standard procedure [30] to obtain Gallager’s bound.
The last step is the optimization of Gallager’s reliability function. Elementary calculus can be used to obtain the closed form of the reliability function.
B-B Decomposition
For a subset $\mathcal{S}\in\mathfrak{S}_{k}^{N}$ of size $k$, we decompose a $k$-sparse vector $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})$ in (12) into $\boldsymbol{x}_{\mathcal{S}}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ for $\boldsymbol{m}_{1}\in\mathcal{M}^{k-w}$ and $\boldsymbol{m}_{2}\in\mathcal{M}^{w}$ satisfying $\boldsymbol{m}=(\boldsymbol{m}_{1},\boldsymbol{m}_{2})$, with
$\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}=\left\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k-w}^{k}(\mathcal{U}):\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}^{\prime}}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}_{1}}\right\},$ (53)
$\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}=\left\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{w}^{N-k}(\mathcal{U}):\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}^{\prime}}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}_{2}}\right\}.$ (54)
The former type $\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ contains all possible candidates of sparse vectors for the support set $\mathcal{S}$ while the latter type $\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ includes them for the remaining (Nk)(N-k) elements.
Fix $w\in\{0,\ldots,k\}$, $w^{\prime}\in\{0,\ldots,k-w\}$, $\boldsymbol{m}\in\mathcal{M}^{k}$, and $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$. If $\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ and $N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}$ hold, there is some $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ such that $p(\boldsymbol{y}|\boldsymbol{x}^{\prime})\geq p(\boldsymbol{y}|\boldsymbol{x})$ and $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}=w^{\prime}$ are satisfied. Decompose $\boldsymbol{x}^{\prime}$ into the two vectors $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{x}^{\prime}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}^{\prime}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$. Since $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}$ is a deterministic function of $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{x}_{1}^{\prime}$, we rewrite $N_{\boldsymbol{x},\boldsymbol{x}^{\prime}}$ as $N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}$.
We evaluate $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$, in which the distribution is induced from that of $\boldsymbol{y}$ conditioned on $\boldsymbol{x}$ in (1). Using the union bound with respect to $\boldsymbol{x}_{1}^{\prime}$ yields
$\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}1(N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}=w^{\prime})P_{\boldsymbol{x}}(\boldsymbol{x}_{1}^{\prime}),$ (55)
with
$P_{\boldsymbol{x}}(\boldsymbol{x}_{1}^{\prime})=\mathbb{P}\left(\left.\bigcup_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}\left\{\frac{p(\boldsymbol{y}|\boldsymbol{x}^{\prime})}{p(\boldsymbol{y}|\boldsymbol{x})}\geq 1\right\}\right|\boldsymbol{x}\right).$ (56)
B-C Gallager’s Bound
We follow Gallager [30] to evaluate $P_{\boldsymbol{x}}(\boldsymbol{x}_{1}^{\prime})$ as
$P_{\boldsymbol{x}}(\boldsymbol{x}_{1}^{\prime})=\mathbb{E}\left[\left.\mathbb{P}\left(\left.\bigcup_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}\left\{\frac{p(\boldsymbol{y}|\boldsymbol{x}^{\prime})}{p(\boldsymbol{y}|\boldsymbol{x})}\geq 1\right\}\right|\boldsymbol{x},\boldsymbol{y}\right)\right|\boldsymbol{x}\right]$
$\leq\mathbb{E}\left[\left.\left\{\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}\mathbb{P}\left(\left.\frac{p(\boldsymbol{y}|\boldsymbol{x}^{\prime})}{p(\boldsymbol{y}|\boldsymbol{x})}\geq 1\right|\boldsymbol{x},\boldsymbol{y}\right)\right\}^{\rho}\right|\boldsymbol{x}\right]$ (57)
for all $\rho\in[0,1]$. The upper bound (57) is trivial when the summation inside the curly brackets $\{\cdots\}$ is larger than $1$. Otherwise, it follows from the inequality $a\leq a^{\rho}$ for all $a\in[0,1]$ and $\rho\in[0,1]$, as well as the union bound. In particular, the upper bound (57) is equivalent to the union bound for $\rho=1$. Using the representation $\mathbb{P}(p(\boldsymbol{y}|\boldsymbol{x}^{\prime})/p(\boldsymbol{y}|\boldsymbol{x})\geq 1|\boldsymbol{x},\boldsymbol{y})=1(p(\boldsymbol{y}|\boldsymbol{x}^{\prime})\geq p(\boldsymbol{y}|\boldsymbol{x}))$ and the inequality $1(p^{\prime}\geq p)\leq(p^{\prime}/p)^{s}$ for all $s\geq 0$, $p>0$, and $p^{\prime}\geq 0$, we obtain
$P_{\boldsymbol{x}}(\boldsymbol{x}_{1}^{\prime})\leq\mathbb{E}\left[\left.\left\{\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}\left[\frac{p(\boldsymbol{y}|\boldsymbol{x}^{\prime})}{p(\boldsymbol{y}|\boldsymbol{x})}\right]^{s}\right\}^{\rho}\right|\boldsymbol{x}\right]$ (58)
for all $\rho\in[0,1]$ and $s\geq 0$.
To evaluate the upper bound (58), we focus on each term in the summation. Using the definition of the AWGN channel (1), as evaluated in [30, Chapter 7], we obtain
$\left[\frac{p(\boldsymbol{y}|\boldsymbol{x}^{\prime})}{p(\boldsymbol{y}|\boldsymbol{x})}\right]^{s}=\exp\left(\frac{s\|\boldsymbol{\omega}\|_{2}^{2}}{2\sigma_{N/k}^{2}}-\frac{s\|\boldsymbol{\omega}+\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{2}^{2}}{2\sigma_{N/k}^{2}}\right)$
$=\exp\left\{-\frac{s(\|\boldsymbol{x}_{1}-\boldsymbol{x}_{1}^{\prime}\|_{2}^{2}+\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2})}{2\sigma_{N/k}^{2}}+\frac{s\boldsymbol{\omega}^{\mathrm{T}}(\boldsymbol{x}^{\prime}-\boldsymbol{x})}{\sigma_{N/k}^{2}}\right\},$ (59)
with $\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}=0$ for $w=0$, where the last equality is due to $\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{2}^{2}=\|\boldsymbol{x}_{1}-\boldsymbol{x}_{1}^{\prime}\|_{2}^{2}+\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}$ with $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$, obtained from the definitions of $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ in (53) and (54). Substituting (59) into the upper bound (58) and using the representation $\boldsymbol{\omega}^{\mathrm{T}}(\boldsymbol{x}^{\prime}-\boldsymbol{x})=\boldsymbol{\omega}_{1}^{\mathrm{T}}(\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1})+\boldsymbol{\omega}_{2}^{\mathrm{T}}\boldsymbol{x}_{2}^{\prime}$, with $\boldsymbol{\omega}_{1}=\boldsymbol{\omega}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{\omega}_{2}=\boldsymbol{\omega}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$, we arrive at
$\exp\left\{\frac{\rho s(\|\boldsymbol{x}_{1}-\boldsymbol{x}_{1}^{\prime}\|_{2}^{2}+\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2})}{2\sigma_{N/k}^{2}}\right\}P_{\boldsymbol{x}}(\boldsymbol{x}_{1}^{\prime})$
$\leq\mathbb{E}\left[\left.e^{\frac{\rho s\boldsymbol{\omega}_{1}^{\mathrm{T}}(\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1})}{\sigma_{N/k}^{2}}}\right|\boldsymbol{x}\right]\mathbb{E}\left[\left.\left\{\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}e^{\frac{s\boldsymbol{\omega}_{2}^{\mathrm{T}}\boldsymbol{x}_{2}^{\prime}}{\sigma_{N/k}^{2}}}\right\}^{\rho}\right|\boldsymbol{x}\right]$
$\leq e^{\frac{\rho^{2}s^{2}\|\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1}\|_{2}^{2}}{2\sigma_{N/k}^{2}}}\left\{\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}e^{s^{2}\|\boldsymbol{x}_{2}^{\prime}\|_{2}^{2}/(2\sigma_{N/k}^{2})}\right\}^{\rho}$
$=\binom{N-k}{w}^{\rho}\exp\left(\frac{\rho^{2}s^{2}\|\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1}\|_{2}^{2}+\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}\rho s^{2}}{2\sigma_{N/k}^{2}}\right).$ (60)
In the derivation of the second inequality, we have used Jensen’s inequality $\mathbb{E}[\{\cdots\}^{\rho}|\boldsymbol{x}]\leq\{\mathbb{E}[\cdots|\boldsymbol{x}]\}^{\rho}$ for all $\rho\in[0,1]$, as well as the fact that $\boldsymbol{\omega}_{1}^{\mathrm{T}}(\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1})/\sigma_{N/k}$ and $\boldsymbol{\omega}_{2}^{\mathrm{T}}\boldsymbol{x}_{2}^{\prime}/\sigma_{N/k}$ conditioned on $\boldsymbol{x}$ follow the zero-mean Gaussian distribution with variance $\|\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1}\|_{2}^{2}$ and $\|\boldsymbol{x}_{2}^{\prime}\|_{2}^{2}$, respectively. The last equality follows from the definition of $\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ in (54) and $\|\boldsymbol{x}_{2}^{\prime}\|_{2}^{2}=\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}$ for all $\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$.
We are ready to evaluate $\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$ in (55). Applying the upper bound (60) to (55), we have
$\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\binom{N-k}{w}^{\rho}\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}1(N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}=w^{\prime})(N/k)^{\frac{\rho f(s,\rho)}{2\sigma^{2}}},$ (61)
where $f(s,\rho)$ is given by
$f(s,\rho)=(a\rho+b)s^{2}-(a+b)s,$ (62)
with $a=\|\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1}\|_{2}^{2}$ and $b=\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}$. Minimizing $f(s,\rho)$ over $s\geq 0$ for fixed $\rho\in[0,1]$ yields $f(s,\rho)\geq-(a+b)^{2}/\{4(a\rho+b)\}$ at $s=(a+b)/\{2(a\rho+b)\}$. Using this lower bound and the upper bound $\binom{N-k}{w}\leq\binom{N}{w}$, we obtain
$\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\binom{N}{w}^{\rho}\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}1(N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}=w^{\prime})(N/k)^{-\frac{(a+b)^{2}\rho}{8\sigma^{2}(a\rho+b)}}.$ (63)
For $w=0$ we recall $b=\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}=0$ to find
$\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{0,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq 1(N_{\boldsymbol{x}_{1},\boldsymbol{u}_{\boldsymbol{m}}}=w^{\prime})(N/k)^{-\frac{\|\boldsymbol{u}_{\boldsymbol{m}}-\boldsymbol{x}_{1}\|_{2}^{2}}{8\sigma^{2}}}.$ (64)
Let $d_{\mathrm{min}}=\min_{m,m^{\prime}\in\mathcal{M}:m\neq m^{\prime}}|u_{m}-u_{m^{\prime}}|>0$ denote the minimum distance between different discrete points in $\mathcal{U}$. The condition $N_{\boldsymbol{x}_{1},\boldsymbol{u}_{\boldsymbol{m}}}=w^{\prime}$ holds only when the elements of $\boldsymbol{u}_{\boldsymbol{m}}$ are different from those of $\boldsymbol{x}_{1}$ in $w^{\prime}$ positions. Thus, we have the lower bound $\|\boldsymbol{u}_{\boldsymbol{m}}-\boldsymbol{x}_{1}\|_{2}^{2}\geq d_{\mathrm{min}}^{2}w^{\prime}$. Evaluating the summation over $\boldsymbol{m}\in\mathcal{M}^{k}$ yields
$\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{0,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\binom{k}{w^{\prime}}(M-1)^{w^{\prime}}(N/k)^{-\frac{d_{\mathrm{min}}^{2}w^{\prime}}{8\sigma^{2}}},$ (65)
which is equivalent to (15) for $w=0$.
For $w>0$, on the other hand, we use the upper bound $\binom{n}{m}\leq(en/m)^{m}$ for (63) to obtain
$\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq(ek/w)^{w}\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}1(N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}=w^{\prime})(N/k)^{-E(\sigma^{2};\rho,a,b)}$ (66)
for all $w\in\{1,\ldots,k\}$ and $\rho\in[0,1]$, with
$E(\sigma^{2};\rho,a,b)=\frac{(a+b)^{2}\rho}{8\sigma^{2}(a\rho+b)}-\rho w,$ (67)
where the last inequality is due to $w\leq k$ and $\rho\leq 1$.
B-D Reliability Function
For $w>0$, let $a=\|\boldsymbol{x}_{1}^{\prime}-\boldsymbol{x}_{1}\|_{2}^{2}$ and $b=\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}$. For all $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$, $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$, and $\boldsymbol{m}\in\mathcal{M}^{k}$ under the condition $N_{\boldsymbol{x}_{1},\boldsymbol{x}_{1}^{\prime}}=w^{\prime}$, we have $a\geq u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$ and $b\geq u_{\mathrm{min}}^{2}w$. Define the reliability function as
$E_{w,w^{\prime}}(\sigma^{2})=\min_{a\geq u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime},b\geq u_{\mathrm{min}}^{2}w}\max_{\rho\in[0,1]}E(\sigma^{2};\rho,a,b).$ (68)
Applying this definition to the upper bound (66) yields
$\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq(ek/w)^{w}M^{w}\binom{k-w}{w^{\prime}}(M-1)^{w^{\prime}}(N/k)^{-E_{w,w^{\prime}}(\sigma^{2})}.$ (69)
To complete the proof of Lemma 1, it is sufficient to prove that $E_{w,w^{\prime}}(\sigma^{2})$ reduces to (17).
We evaluate $E_{w,w^{\prime}}(\sigma^{2})$. Let $\tilde{a}=(a+b)^{2}/(8\sigma^{2}a)$ and $\tilde{b}=b/a$. The function $E(\sigma^{2};\rho,a,b)$ is represented as
$E(\sigma^{2};\rho,a,b)=\frac{\tilde{a}\rho}{\rho+\tilde{b}}-w\rho.$ (70)
It is straightforward to confirm the concavity of $E$ with respect to $\rho\in[0,1]$. Thus, $E$ is maximized with respect to $\rho$ at its stationary point, $\rho=0$, or $\rho=1$. Evaluating the derivative of $E$ yields
$\frac{\partial E}{\partial\rho}=\frac{\tilde{a}}{\rho+\tilde{b}}-\frac{\tilde{a}\rho}{(\rho+\tilde{b})^{2}}-w=-\frac{w(\rho+\tilde{b})^{2}-\tilde{a}\tilde{b}}{(\rho+\tilde{b})^{2}}.$ (71)
The condition $\partial E/\partial\rho\leq 0$ for all $\rho\in[0,1]$ is equivalent to $\tilde{a}\leq w\tilde{b}$. In this case, we find $E\leq E(\sigma^{2};0,a,b)=0$. Furthermore, $\tilde{a}\leq w\tilde{b}$ is equivalent to $8\sigma^{2}w\geq(a+b)^{2}/b$. Minimizing the right-hand side (RHS) over $a\geq u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$ and $b\geq u_{\mathrm{min}}^{2}w$, we obtain
$\frac{(a+b)^{2}}{b}$ $\geq\left(\frac{u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}}{\sqrt{b}}+\sqrt{b}\right)^{2}$
$\geq 4(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}).$ (72)
Here, the equality for the first inequality holds at $a=u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$. The second inequality is due to the arithmetic-mean and geometric-mean (AM-GM) inequality, in which the equality holds at $b=u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$. Thus, we find $E_{w,w^{\prime}}(\sigma^{2})=0$ for $\sigma^{2}\geq(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})/(2w)$.
The condition $\partial E/\partial\rho>0$ for all $\rho\in[0,1]$ is equivalent to $(\tilde{a}\tilde{b}/w)^{1/2}>\tilde{b}+1$. In this case we have
$E\leq E(\sigma^{2};1,a,b)=\frac{\tilde{a}}{1+\tilde{b}}-w=\frac{a+b}{8\sigma^{2}}-w.$ (73)
Minimizing the upper bound over $a\geq u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$ and $b\geq u_{\mathrm{min}}^{2}w$, we have $E\leq(2u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})/(8\sigma^{2})-w$ at $a=u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$ and $b=u_{\mathrm{min}}^{2}w$. Furthermore, $(\tilde{a}\tilde{b}/w)^{1/2}>\tilde{b}+1$ is equivalent to $b>8\sigma^{2}w$, which reduces to $\sigma^{2}<u_{\mathrm{min}}^{2}/8$ for the minimizer $b=u_{\mathrm{min}}^{2}w$.
Finally, assume $\tilde{a}>w\tilde{b}$ and $(\tilde{a}\tilde{b}/w)^{1/2}\leq\tilde{b}+1$. In this case, $E$ is maximized at $\rho=(\tilde{a}\tilde{b}/w)^{1/2}-\tilde{b}\equiv\rho_{\mathrm{s}}$ satisfying $\partial E/\partial\rho=0$ for $\rho=\rho_{\mathrm{s}}$ to obtain
$E\leq E(\sigma^{2};\rho_{\mathrm{s}},a,b)=\{\tilde{a}^{1/2}-(w\tilde{b})^{1/2}\}^{2}$
$=\frac{\{a+(\sqrt{b}-\sqrt{2\sigma^{2}w})^{2}-2\sigma^{2}w\}^{2}}{8\sigma^{2}a}\equiv E_{*}.$ (74)
Furthermore, $\tilde{a}>w\tilde{b}$ and $(\tilde{a}\tilde{b}/w)^{1/2}\leq\tilde{b}+1$ are equivalent to $8\sigma^{2}w\in[b,(a+b)^{2}/b)$.
We minimize the upper bound $E_{*}$ over $a\geq u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$ and $b\geq u_{\mathrm{min}}^{2}w$. Since we have already evaluated $E_{w,w^{\prime}}(\sigma^{2})$ for $\sigma^{2}\geq(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})/(2w)$ and $\sigma^{2}<u_{\mathrm{min}}^{2}/8$, we can assume $\sigma^{2}\in[u_{\mathrm{min}}^{2}/8,(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})/(2w))$, which implies $a\geq u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}>2\sigma^{2}w$. For $\sigma^{2}\in[u_{\mathrm{min}}^{2}/2,(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})/(2w))$, we use $b\geq u_{\mathrm{min}}^{2}w$ and $2\sigma^{2}w\geq u_{\mathrm{min}}^{2}w$ to have
$E_{*}\geq\frac{(a-2\sigma^{2}w)^{2}}{8\sigma^{2}a},$ (75)
where the equality holds at $b=2\sigma^{2}w$. The lower bound is monotonically increasing with respect to $a\geq 2\sigma^{2}w$. Thus, the minimum is attained at $a=u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}\geq 2\sigma^{2}w$,
$E_{*}\geq\frac{(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}-2\sigma^{2}w)^{2}}{8\sigma^{2}(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})}.$ (76)
For $\sigma^{2}\in[u_{\mathrm{min}}^{2}/8,u_{\mathrm{min}}^{2}/2)$, we find $2\sigma^{2}w<u_{\mathrm{min}}^{2}w\leq b$ to lower-bound $E_{*}$ in (74) as
$E_{*}\geq\frac{\{a-(2\sqrt{2\sigma^{2}u_{\mathrm{min}}^{2}}-u_{\mathrm{min}}^{2})w\}^{2}}{8\sigma^{2}a}$ (77)
where the equality holds at $b=u_{\mathrm{min}}^{2}w$. The lower bound is monotonically increasing with respect to $a\geq(2\sqrt{2\sigma^{2}u_{\mathrm{min}}^{2}}-u_{\mathrm{min}}^{2})w\geq 0$ for $\sigma^{2}\geq u_{\mathrm{min}}^{2}/8$. Since the assumption $\sigma^{2}<u_{\mathrm{min}}^{2}/2$ implies $(2\sqrt{2\sigma^{2}u_{\mathrm{min}}^{2}}-u_{\mathrm{min}}^{2})w<u_{\mathrm{min}}^{2}w$, the minimum is attained at $a=u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime}$. Thus, we arrive at
$E_{*}\geq\frac{\{2(u_{\mathrm{min}}^{2}-\sqrt{2\sigma^{2}u_{\mathrm{min}}^{2}})w+d_{\mathrm{min}}^{2}w^{\prime}\}^{2}}{8\sigma^{2}(u_{\mathrm{min}}^{2}w+d_{\mathrm{min}}^{2}w^{\prime})}$ (78)
for $\sigma^{2}\in[u_{\mathrm{min}}^{2}/8,u_{\mathrm{min}}^{2}/2)$. Thus, $E_{w,w^{\prime}}(\sigma^{2})$ reduces to (17).
Appendix C Evaluation of (14) for $M>1$
Assume $M>1$. For $w=0$ we use Lemma 1, M1<MM-1<M, and $\binom{k}{w^{\prime}}\leq(ek/w^{\prime})^{w^{\prime}}$ to obtain
$\sum_{w^{\prime}=d}^{k}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{0,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\sum_{w^{\prime}=d}^{k}(ek/d)^{w^{\prime}}M^{w^{\prime}}(N/k)^{-\frac{d_{\mathrm{min}}^{2}w^{\prime}}{8\sigma^{2}}}\to 0,$ (79)
where the last convergence follows from Proposition 2 with $r_{k,N}=(ek/d)M(N/k)^{-d_{\mathrm{min}}^{2}/(8\sigma^{2})}$.
For $w>0$ we use Lemma 1 and the trivial upper bound M1<MM-1<M to have
$\sum_{w+w^{\prime}\geq d}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\sum_{w+w^{\prime}\geq d}(ek/w)^{w}\binom{k-w}{w^{\prime}}M^{w+w^{\prime}}(N/k)^{-E_{w,w^{\prime}}(\sigma^{2})}.$ (80)
In particular, for $w^{\prime}=0$ we use the linearity of $E_{w,0}(\sigma^{2})$ with respect to $w$ to obtain
$\sum_{w=d}^{k}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=0|\boldsymbol{x})$
$\leq\sum_{w=d}^{k}(ek/d)^{w}M^{w}(N/k)^{-wE_{1,0}(\sigma^{2})}\to 0,$ (81)
where the last convergence follows from Proposition 2 with $r_{k,N}=(ek/d)M(N/k)^{-E_{1,0}(\sigma^{2})}$.
To evaluate the upper bound (80) for $w>0$ and $w^{\prime}>0$, we prove two technical results.
Proposition 3
Let $\mathcal{W}_{i}=\{(w,w^{\prime}):w\in\{1,\ldots,k\},w^{\prime}\in\{1,\ldots,k-w\},w+w^{\prime}=i\}$ for $i\in\{2,\ldots,k\}$. Then,
$\min_{(w,w^{\prime})\in\mathcal{W}_{i}}w^{w}w^{\prime w^{\prime}}\geq(i/2)^{i}.$ (82)
Proof:
Define $f(w,w^{\prime})=\log(w^{w}w^{\prime w^{\prime}})=w\log w+w^{\prime}\log w^{\prime}$. The minimizer of $f(w,w^{\prime})$ over $(w,w^{\prime})\in\mathcal{W}_{i}$ is also the minimizer of the original minimization problem. Relax the feasible region $\mathcal{W}_{i}$ to the linear constraints $\mathcal{W}_{i}^{\mathrm{L}}=\{(w,w^{\prime})\in\mathbb{R}^{2}:w\in[1,k],w^{\prime}\in[1,k-w],w+w^{\prime}=i\}$. Since $f(w,w^{\prime})$ is convex on $(w,w^{\prime})\in\mathcal{W}_{i}^{\mathrm{L}}$, the minimization of $f$ over $\mathcal{W}_{i}^{\mathrm{L}}$ is convex programming.
We substitute $w^{\prime}=i-w$ into $f(w,w^{\prime})$ to obtain $f(w,i-w)=w\log w+(i-w)\log(i-w)\equiv f(w)$ for $w\in[1,i-1]$. Differentiating $f(w)$ yields
$f^{\prime}(w)=\log w-\log(i-w),$ (83)
which implies $f^{\prime}(w)=0$ only at $w=i/2$. Thus, $f(w,w^{\prime})$ is bounded from below by $f(i/2)$, which is equivalent to the lower bound (82). ∎
Proposition 4
Let $a=2(u_{\mathrm{min}}^{2}-\sqrt{2\sigma^{2}u_{\mathrm{min}}^{2}})/u_{\mathrm{min}}^{2}$. Then,
$8\sigma^{2}E_{w,w^{\prime}}^{(2)}(\sigma^{2})\geq a^{2}u_{\mathrm{min}}^{2}w+ad_{\mathrm{min}}^{2}w^{\prime}$ (84)
holds for all $w>0$, $w^{\prime}>0$, and $\sigma^{2}\in[u_{\mathrm{min}}^{2}/8,u_{\mathrm{min}}^{2}/2)$.
Proof:
Let $\tilde{w}=u_{\mathrm{min}}^{2}w$ and $\tilde{w}^{\prime}=d_{\mathrm{min}}^{2}w^{\prime}$. We rewrite $E_{w,w^{\prime}}^{(2)}(\sigma^{2})$ in (19) as
$E_{w,w^{\prime}}^{(2)}(\sigma^{2})=\frac{(a\tilde{w}+\tilde{w}^{\prime})^{2}}{8\sigma^{2}(\tilde{w}+\tilde{w}^{\prime})}$. (85)
Since $a\in(0,1]$ holds for all $\sigma^{2}\in[u_{\mathrm{min}}^{2}/8,u_{\mathrm{min}}^{2}/2)$, we have $\tilde{w}^{\prime}\geq a\tilde{w}^{\prime}$, which is equivalent to
$\frac{a\tilde{w}+\tilde{w}^{\prime}}{\tilde{w}+\tilde{w}^{\prime}}\geq a.$ (86)
Multiplying both sides by $a\tilde{w}+\tilde{w}^{\prime}$, we obtain (84). ∎
We evaluate the upper bound (80) for $w>0$ and $w^{\prime}>0$. Let $\mathcal{W}=\{(w,w^{\prime}):w\in\{1,\ldots,k\},w^{\prime}\in\{1,\ldots,k-w\},w+w^{\prime}\geq d\}$. By definition, we have $\mathcal{W}=\cup_{i=d}^{k}\mathcal{W}_{i}$. Furthermore, define $E_{\mathrm{min}}=(8\sigma^{2})^{-1}\min\{a^{2}u_{\mathrm{min}}^{2},ad_{\mathrm{min}}^{2}\}>0$. Using the upper bound $\binom{k-w}{w^{\prime}}\leq(ek/w^{\prime})^{w^{\prime}}$ and the lower bounds (82) and (84) yields
$\sum_{(w,w^{\prime})\in\mathcal{W}}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\mathbb{P}(\hat{\boldsymbol{x}}_{\mathrm{ML}}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}}),N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}=w^{\prime}|\boldsymbol{x})$
$\leq\sum_{i=d}^{k}(i/2)^{-i}\sum_{(w,w^{\prime})\in\mathcal{W}_{i}}(ekM)^{w+w^{\prime}}(N/k)^{-E_{\mathrm{min}}(w+w^{\prime})}$
$\leq\sum_{i=d}^{k}(i-1)r_{k,N}^{i},$ (87)
with $r_{k,N}=(2ek/d)M(N/k)^{-E_{\mathrm{min}}}$, where the last inequality follows from the upper bound $(2ekM/i)^{i}(N/k)^{-E_{\mathrm{min}}i}\leq r_{k,N}^{i}$ for all $i\in\{d,\ldots,k\}$.
We prove that the upper bound (87) converges to zero in the sublinear sparsity limit. Consider the following geometric series:
$S(r)=\sum_{i=d}^{k}r^{i}=r^{d}\frac{1-r^{k-d+1}}{1-r}.$ (88)
Using $S^{\prime}(r)=\sum_{i=d}^{k}ir^{i-1}$, we find $\sum_{i=d}^{k}ir^{i}=rS^{\prime}(r)$. We evaluate the derivative of $S(r)$ at $r=r_{k,N}$ to obtain
$\sum_{i=d}^{k}ir_{k,N}^{i}=\frac{dr_{k,N}^{d}-(k+1)r_{k,N}^{k+1}}{1-r_{k,N}}+\frac{r_{k,N}^{d+1}-r_{k,N}^{k+2}}{(1-r_{k,N})^{2}}\to 0$ (89)
in the sublinear sparsity limit, where the last convergence follows from the definitions $d=\lceil k/\sqrt{\log(N/k)}\rceil$ and $r_{k,N}=(2ek/d)M(N/k)^{-E_{\mathrm{min}}}$. This implies the convergence of the upper bound (87) to zero in the sublinear sparsity limit. Combining these results and the upper bound (14), we arrive at $\mathbb{P}(\mathcal{E}_{d}|\boldsymbol{x})\to 0$ in the sublinear sparsity limit for $M>1$.
Appendix D Proof of Lemma 2
D-A Overview
Without loss of generality, we assume that $\min_{m\in\mathcal{M}}|u_{m}|$ is attained at positive $u_{m}>0$. The proof of Lemma 2 consists of three steps: A first step is the formulation of the KL divergence $D(p\|q)$. To obtain a tight upper bound on the KL divergence, we utilize Jensen’s inequality only for Gaussian noise samples in positions where the signal vector $\boldsymbol{x}$ has zeros. If we applied Jensen’s inequality to all Gaussian noise samples, we would have a loose upper bound on the KL divergence.
In a second step, a truncation technique is used for the remaining $k$ Gaussian noise samples. The threshold in the truncation separates the Gaussian noise samples into typical samples and non-typical samples. The set of typical samples contains all $k$ Gaussian noise samples that occur jointly with high probability, while the set of non-typical samples consists of the remaining rare samples. In the second step, the upper bound in the first step is evaluated for the set of typical samples. This evaluation itself is not challenging because typical noise samples are bounded.
The last step is the technically most challenging part in the proof of Lemma 2—evaluation for non-typical noise samples. Let $w\in\{1,\ldots,k\}$ denote the number of incorrectly detected positions, as considered in the direct part. The main challenge is to upper-bound a Gaussian mixture associated with $k$-dimensional and (kw)(k-w)-sparse vectors by a single Gaussian pdf for a dense vector. This upper bound is proved via a graphical technique based on Hall’s marriage theorem [41]. It is enough tight to evaluate the upper bound in the first step for non-typical noise samples.
D-B KL Divergence
We recall that the true marginal pdf of $\boldsymbol{y}$ is induced from the conditional pdf $p(\boldsymbol{y}|\boldsymbol{x})$ defined via (1) and the true distribution of $\boldsymbol{x}$—uniform distribution on $\mathcal{X}_{k}^{N}(\mathcal{U})$. Let
$\lambda=1+\frac{\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}{N\sigma_{N/k}^{2}}.$ (90)
From the definition of $\boldsymbol{y}$ in (1) and the independence between $\boldsymbol{x}$ and $\boldsymbol{y}$, the true expectation $\mathbb{E}_{p}[\|\boldsymbol{y}\|_{2}^{2}]$ is equal to $\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]+N\sigma_{N/k}^{2}=N\lambda\sigma_{N/k}^{2}$. By definition, $q(\boldsymbol{y})$ is represented as
$q(\boldsymbol{y})=\frac{1}{(2\pi\lambda\sigma_{N/k}^{2})^{N/2}}\exp\left(-\frac{\|\boldsymbol{y}\|_{2}^{2}}{2\lambda\sigma_{N/k}^{2}}\right).$ (91)
To evaluate the KL divergence $D(p\|q)$, we first evaluate $\log\{p(\boldsymbol{y})/q(\boldsymbol{y})\}$. We use the definition of $p(\boldsymbol{y}|\boldsymbol{x})$ in (1) to obtain
$p(\boldsymbol{y})=\frac{e^{-\|\boldsymbol{y}\|_{2}^{2}/(2\sigma_{N/k}^{2})}}{(2\pi\sigma_{N/k}^{2})^{N/2}|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})}e^{(\boldsymbol{x}^{\mathrm{T}}\boldsymbol{y}-\|\boldsymbol{x}\|_{2}^{2}/2)/\sigma_{N/k}^{2}}.$ (92)
Combining the two results (91) and (92), we arrive at
$\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}=\frac{N}{2}\log\lambda-\frac{(\lambda-1)\|\boldsymbol{y}\|_{2}^{2}}{2\lambda\sigma_{N/k}^{2}}$
$+\log\left\{\frac{1}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})}e^{(\boldsymbol{x}^{\mathrm{T}}\boldsymbol{y}-\|\boldsymbol{x}\|_{2}^{2}/2)/\sigma_{N/k}^{2}}\right\}.$ (93)
We upper-bound the KL divergence $D(p\|q)$. Evaluating the expectation of (93) over $p(\boldsymbol{y})$ and subsequently using the definition of $\lambda$ in (90) and the upper bound $\log(1+x)\leq x$ for all $x\geq 0$, we have
$D(p\|q)\leq\frac{1}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})}g(\boldsymbol{x}),$ (94)
with
$g(\boldsymbol{x})=\mathbb{E}\left[\log\left\{\frac{1}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{U})}e^{\frac{\boldsymbol{x}^{\prime\mathrm{T}}(\boldsymbol{x}+\boldsymbol{\omega})-\|\boldsymbol{x}^{\prime}\|_{2}^{2}/2}{\sigma_{N/k}^{2}}}\right\}\right].$ (95)
To derive a tight upper bound on $g(\boldsymbol{x})$ for all $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$, we use the representation $\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ of $\mathcal{X}_{k}^{N}(\mathcal{U})$ in (13) to obtain
$g(\boldsymbol{x})=\mathbb{E}\left[\log\left\{\frac{1}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{w=0}^{k}g_{k-w}(\boldsymbol{x},\boldsymbol{\omega})\right\}\right],$ (96)
with
$g_{k-w}(\boldsymbol{x},\boldsymbol{\omega})=\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}e^{-\frac{\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}}{2\sigma_{N/k}^{2}}}\sum_{\boldsymbol{x}^{\prime}\in\mathcal{T}_{k,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})}e^{\frac{\boldsymbol{x}^{\prime\mathrm{T}}(\boldsymbol{x}+\boldsymbol{\omega})}{\sigma_{N/k}^{2}}}.$ (97)
Let $|\boldsymbol{x}|$ denote the element-wise absolute values of $\boldsymbol{x}$, i.e. $|\boldsymbol{x}|]_{n}=|[\boldsymbol{x}]_{n}|$. Decomposing $\boldsymbol{x}^{\prime}\in\mathcal{T}_{k,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ into $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ in (53) and (54) and subsequently using the upper bound $\boldsymbol{x}_{1}^{\prime\mathrm{T}}\boldsymbol{x}_{1}\leq|\boldsymbol{x}_{1}^{\prime}|^{\mathrm{T}}|\boldsymbol{x}_{1}|$ with $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$, we have
$g_{k-w}(\boldsymbol{x},\boldsymbol{\omega})\leq\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}e^{-\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}/(2\sigma_{N/k}^{2})}$
$\cdot\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}e^{(|\boldsymbol{x}_{1}^{\prime}|^{\mathrm{T}}|\boldsymbol{x}_{1}|+\boldsymbol{x}_{1}^{\prime\mathrm{T}}\boldsymbol{\omega}_{1})/\sigma_{N/k}^{2}}\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}e^{\boldsymbol{x}_{2}^{\prime\mathrm{T}}\boldsymbol{\omega}_{2}/\sigma_{N/k}^{2}},$ (98)
with $\boldsymbol{\omega}_{1}=\boldsymbol{\omega}_{\mathcal{S}_{\boldsymbol{x}}}$ and $\boldsymbol{\omega}_{2}=\boldsymbol{\omega}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{N-k})$.
We use Jensen’s inequality with respect to $\boldsymbol{\omega}_{2}$ conditioned on $\boldsymbol{\omega}_{1}$ to evaluate $g(\boldsymbol{x})$ in (96) as
$g(\boldsymbol{x})\leq\mathbb{E}\left[\log\left\{\sum_{w=0}^{k}\frac{|\mathcal{M}|^{w}}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\binom{N-k}{w}G_{k-w}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})\right\}\right],$ (99)
with $G_{0}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})=1$ and
$G_{k-w}^{k}$ $=\frac{1}{|\mathcal{M}|^{w}}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}e^{\frac{|\boldsymbol{x}_{1}^{\prime}|^{\mathrm{T}}|\boldsymbol{x}_{1}|-2^{-1}\|\boldsymbol{x}_{1}^{\prime}\|_{2}^{2}+\boldsymbol{x}_{1}^{\prime\mathrm{T}}\boldsymbol{\omega}_{1}}{\sigma_{N/k}^{2}}}$
$=\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{X}_{k-w}^{k}(\mathcal{U})}e^{\frac{|\boldsymbol{x}_{1}^{\prime}|^{\mathrm{T}}|\boldsymbol{x}_{1}|-2^{-1}\|\boldsymbol{x}_{1}^{\prime}\|_{2}^{2}+\boldsymbol{x}_{1}^{\prime\mathrm{T}}\boldsymbol{\omega}_{1}}{\sigma_{N/k}^{2}}}$ (100)
for $w\in\{0,\ldots,k-1\}$, where the last equality follows from the definition of $\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ in (53).
We derive a uniform upper bound on $G_{k-w}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})$ with respect to $\boldsymbol{x}_{1}$. Represent (100) as
$G_{k-w}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})=\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{X}_{k-w}^{k}(\{1\})}\prod_{j\in\mathcal{S}_{\boldsymbol{x}_{1}^{\prime}}}h([\boldsymbol{x}_{1}]_{j},[\boldsymbol{\omega}_{1}]_{j}),$ (101)
with
$h(u,\omega)=\sum_{m\in\mathcal{M}}\exp\left(\frac{|u_{m}||u|}{\sigma_{N/k}^{2}}-\frac{u_{m}^{2}}{2\sigma_{N/k}^{2}}+\frac{u_{m}\omega}{\sigma_{N/k}^{2}}\right).$ (102)
We note that $u$ satisfying $|u|=u_{\mathrm{max}}$ is the maximizer of $h(u,\omega)$ over $u\in\mathcal{U}\cup\{0\}$. Thus, we have the uniform upper bound $G_{k-w}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})\leq G_{k-w}^{k}(\boldsymbol{\omega}_{1})$, with $G_{0}^{k}(\boldsymbol{\omega}_{1})=1$ and
$G_{k-w}^{k}(\boldsymbol{\omega}_{1})=\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{X}_{k-w}^{k}(\mathcal{U})}e^{\frac{u_{\mathrm{max}}\|\boldsymbol{x}_{1}^{\prime}\|_{1}-2^{-1}\|\boldsymbol{x}_{1}^{\prime}\|_{2}^{2}+\boldsymbol{x}_{1}^{\prime\mathrm{T}}\boldsymbol{\omega}_{1}}{\sigma_{N/k}^{2}}}$ (103)
for $w\in\{0,\ldots,k-1\}$.
We are ready to evaluate the upper bound (94) on the KL divergence $D(p\|q)$. Applying the uniform upper bound $G_{k-w}^{k}(\boldsymbol{x}_{1},\boldsymbol{\omega}_{1})\leq G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ to $g(\boldsymbol{x})$ in (99) yields
$D(p\|q)\leq\mathbb{E}\left[\log\left\{\sum_{w=0}^{k}\frac{|\mathcal{M}|^{w}}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\binom{N-k}{w}G_{k-w}^{k}(\boldsymbol{\omega}_{1})\right\}\right].$ (104)
Using $|\mathcal{X}_{k}^{N}(\mathcal{U})|=|\mathcal{M}|^{k}\binom{N}{k}$ and the following upper bound:
$\binom{N}{k}^{-1}\binom{N-k}{w}<\frac{k!}{(N-k)^{k}}\frac{(N-k)^{w}}{w!}<(N/k-1)^{w-k},$ (105)
we arrive at
$D(p\|q)<\mathbb{E}\left[\log\left\{\sum_{w=0}^{k}[|\mathcal{M}|(N/k-1)]^{w-k}G_{k-w}^{k}(\boldsymbol{\omega}_{1})\right\}\right].$ (106)
The signal vector $\boldsymbol{x}$ can depend on the upper bound (106) only through the Gaussian vector $\boldsymbol{\omega}_{1}=\boldsymbol{\omega}_{\mathcal{S}_{\boldsymbol{x}}}$. Since $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$ is independent of $\boldsymbol{x}$, however, (106) does not depend on $\boldsymbol{x}$. Thus, without loss of generality, we assume $\mathcal{S}_{\boldsymbol{x}}=\{1,\ldots,k\}$. In this case, $\boldsymbol{\omega}_{1}=\boldsymbol{\omega}_{\mathcal{S}_{\boldsymbol{x}}}$ has the elements $\{\omega_{j}:j\in\{1,\ldots,k\}\}$.
D-C Typical Noise Samples
To evaluate the expectation in (106) over $\boldsymbol{\omega}_{1}$, we write the elements of $\mathcal{X}_{k-w}^{k}(\mathcal{U})$ for each $w\in\{0,\ldots,k-1\}$ as $\{\boldsymbol{x}_{i}^{w}\in\mathcal{X}_{k-w}^{k}(\mathcal{U}):i\in\{1,\ldots,|\mathcal{X}_{k-w}^{k}(\mathcal{U})|\}$. Define the random variable $\Omega_{i}^{w}=(\boldsymbol{x}_{i}^{w})^{\mathrm{T}}\boldsymbol{\omega}_{1}/\sigma_{N/k}^{2}$. We use $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$ to find $\Omega_{i}^{w}\sim\mathcal{N}(0,\|\boldsymbol{x}_{i}^{w}\|_{2}^{2}/\sigma_{N/k}^{2})$. We focus on the $k$ basis random variables $\boldsymbol{\Omega}_{\mathrm{b}}=\{\Omega_{i}^{k-1}:i\in\{1,\ldots,k\}\}$ for $k$ different $1$-sparse vectors $\boldsymbol{x}_{i}^{k-1}\in\mathcal{X}_{1}^{k}(\mathcal{U})$. We assume that $\boldsymbol{x}_{i}^{k-1}$ has a non-zero element with amplitude $u_{\mathrm{max}}$ in the $i$th element. Since all random variables $\{\Omega_{i}^{w}\}$ are linear combinations of the $k$ basis random variables, there are some deterministic linear functions $f_{i}^{w}:\mathbb{R}^{k}\to\mathbb{R}$ such that $\Omega_{i}^{w}=f_{i}^{w}(\boldsymbol{\Omega}_{\mathrm{b}})$ holds.
Consider $\mathcal{E}_{i}^{w}=\{\Omega_{i}^{w}:|\Omega_{i}^{w}|\leq d_{i}^{w}\}$ for some $d_{i}^{w}\in\mathbb{R}$. In particular, we focus on the basis event $\mathcal{E}_{\mathrm{b}}=\cap_{i=1}^{k}\mathcal{E}_{i}^{k-1}$. Using this basis event $\mathcal{E}_{\mathrm{b}}$, we decompose the upper bound on $D(p\|q)$ in (106) into $D(p\|q)\leq g_{1}+g_{2}$, with
$g_{1}=\mathbb{E}\left[1\left(\mathcal{E}_{\mathrm{b}}\right)\log\left\{\sum_{w=0}^{k}[|\mathcal{M}|(N/k-1)]^{w-k}G_{k-w}^{k}(\boldsymbol{\omega}_{1})\right\}\right],$ (107)
$g_{2}=\mathbb{E}\left[1\left(\mathcal{E}_{\mathrm{b}}^{\mathrm{c}}\right)\log\left\{\sum_{w=0}^{k}[|\mathcal{M}|(N/k-1)]^{w-k}G_{k-w}^{k}(\boldsymbol{\omega}_{1})\right\}\right].$ (108)
We need to prove that both $g_{1}$ and $g_{2}$ converge to zero in the sublinear sparsity limit.
We define $d_{i}^{w}$ so that $\mathcal{E}_{\mathrm{b}}^{\mathrm{c}}$ becomes a rare event. Let
$d_{i}^{w}=\frac{\|\boldsymbol{x}_{i}^{w}\|_{1}}{u_{\mathrm{max}}}d_{*},$ (109)
with
$d_{*}=(1-\epsilon)\log(N/k)-\frac{u_{\mathrm{max}}^{2}}{2\sigma_{N/k}^{2}}$ (110)
for any $\epsilon\in(0,1-u_{\mathrm{max}}^{2}/(2\sigma^{2}))$. In this case, we have $d_{i}^{k-1}=d_{*}$ for $i\in\{1,\ldots,k\}$. It is straightforward to confirm $d_{i}^{w}>0$ in (109) for all $i$ and $w$ or $d_{*}>0$ in (110). Using $\sigma_{N/k}^{2}=\sigma^{2}/\log(N/k)$, we have
$d_{*}=\left(1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}-\epsilon\right)\log(N/k)>0,$ (111)
where the last positivity is due to $\epsilon\in(0,1-u_{\mathrm{max}}^{2}/(2\sigma^{2}))$.
To understand the significance of these definitions, we use the definition $\Omega_{i}^{k-1}=u_{*}[\boldsymbol{\omega}_{1}]_{i}/\sigma_{N/k}^{2}$ with $|u_{*}|=u_{\mathrm{max}}$ for all $i\in\{1,\ldots,k\}$ to represent $\Omega_{i}^{w}=(\boldsymbol{x}_{i}^{w})^{\mathrm{T}}\boldsymbol{\omega}_{1}/\sigma_{N/k}^{2}$ as $\Omega_{i}^{w}=u_{*}^{-1}\sum_{j=1}^{k}[\boldsymbol{x}_{i}^{w}]_{j}\Omega_{j}^{k-1}$. Applying the triangle inequality yields the following upper bound:
$|\Omega_{i}^{w}|\leq\frac{1}{u_{\mathrm{max}}}\sum_{j=1}^{k}|[\boldsymbol{x}_{i}^{w}]_{j}||\Omega_{j}^{k-1}|\leq d_{i}^{w}$ (112)
conditioned on $\mathcal{E}_{\mathrm{b}}$, where the last inequality follows from the conditioning $|\Omega_{i}^{k-1}|\leq d_{i}^{k-1}$ and the definition of $d_{i}^{w}$ in (109). Thus, the basis event $\mathcal{E}_{\mathrm{b}}$ implies the event $\mathcal{E}_{i}^{w}$ for all $i$ and $w$.
Lemma 4
For $C_{0}$ and $C_{1,1}$ given in (26), we have
$\mathbb{P}(|\Omega_{i}^{k-1}|>d_{i}^{k-1})\leq\frac{2e^{C_{0}\epsilon}}{(N/k)^{C_{1,1}}},$ (113)
$\mathbb{P}(\mathcal{E}_{\mathrm{b}})>\left\{1-\frac{2e^{C_{0}\epsilon}}{(N/k)^{C_{1,1}}}\right\}^{k}.$ (114)
Proof:
We first evaluate $\mathbb{P}(|\Omega_{i}^{k-1}|>d_{i}^{k-1})$ with $d_{i}^{k-1}=d_{*}$ in (110). Since $\Omega_{i}^{k-1}$ is a zero-mean Gaussian random variable with variance $u_{\mathrm{max}}^{2}/\sigma_{N/k}^{2}$, we use the representation $\mathbb{P}(|\Omega_{i}^{k-1}|>d_{i}^{k-1})=2Q(\sigma_{N/k}d_{*}/u_{\mathrm{max}})$ and the Chernoff bound $Q(x)\leq e^{-x^{2}/2}$ for $x\geq 0$ to obtain
$\mathbb{P}(|\Omega_{i}^{k-1}|>d_{i}^{k-1})\leq 2e^{-\sigma_{N/k}^{2}d_{*}^{2}/(2u_{\mathrm{max}}^{2})}.$ (115)
Evaluating the exponent with $d_{*}$ in (111) as
$\frac{\sigma_{N/k}^{2}d_{*}^{2}}{2u_{\mathrm{max}}^{2}}>\frac{\sigma^{2}}{2u_{\mathrm{max}}^{2}}\left(1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}\right)^{2}\log(N/k)$
$-\frac{\sigma^{2}}{u_{\mathrm{max}}^{2}}\left(1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}\right)\epsilon=C_{1,1}\log(N/k)-C_{0}\epsilon,$ (116)
with $C_{0}$ and $C_{1,1}$ in (26), we arrive at (113).
We next evaluate $\mathbb{P}(\mathcal{E}_{\mathrm{b}})$. Since $\{\Omega_{i}^{k-1}\}_{i=1}^{k}$ are i.i.d. random variables, we represent $\mathbb{P}(\mathcal{E}_{\mathrm{b}})$ as
$\mathbb{P}(\mathcal{E}_{\mathrm{b}})=\prod_{i=1}^{k}\mathbb{P}(|\Omega_{i}^{k-1}|\leq d_{i}^{k-1})=\left\{\mathbb{P}(|\Omega_{1}^{k-1}|\leq d_{*})\right\}^{k}.$ (117)
Using $\mathbb{P}(|\Omega_{1}^{k-1}|\leq d_{*})=1-\mathbb{P}(|\Omega_{1}^{k-1}|>d_{*})$ and the upper bound (113), we arrive at the lower bound (114). ∎
Lemma 4 implies that $\mathcal{E}_{\mathrm{b}}$ is a typical event under the assumption $\gamma<C_{1,1}/(C_{1,1}+1)$: $\mathbb{P}(\mathcal{E}_{\mathrm{b}})\to 1$ holds in the sublinear sparsity limit. Thus, $g_{1}$ in (107) represents contributions from the typical event $\mathcal{E}_{\mathrm{b}}$ while $g_{2}$ in (108) represents those from the non-typical rare event $\mathcal{E}_{\mathrm{b}}^{\mathrm{c}}$.
It is straightforward to evaluate $g_{1}$ in (107). Since $\mathcal{E}_{\mathrm{b}}$ implies $\mathcal{E}_{i}^{w}$, we use $|\Omega_{i}^{w}|\leq d_{i}^{w}$ to evaluate $G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ in (103) as
$G_{k-w}^{k}(\boldsymbol{\omega}_{1})\leq\sum_{i=1}^{|\mathcal{X}_{k-w}^{k}(\mathcal{U})|}e^{\frac{u_{\mathrm{max}}\|\boldsymbol{x}_{i}^{w}\|_{1}-2^{-1}\|\boldsymbol{x}_{i}^{w}\|_{2}^{2}}{\sigma_{N/k}^{2}}+d_{i}^{w}}$ (118)
conditioned on $\mathcal{E}_{\mathrm{b}}$. Substituting the definition of $d_{i}^{w}$ in (109) and using the lower bound $\|\boldsymbol{x}_{i}^{w}\|_{1}\geq u_{\mathrm{min}}(k-w)$ for the coefficient of $\epsilon\log(N/k)$, we have
$G_{k-w}^{k}(\boldsymbol{\omega}_{1})\leq\binom{k}{w}\sum_{\boldsymbol{m}_{1}\in\mathcal{M}^{k-w}}(N/k)^{f(\boldsymbol{u}_{\boldsymbol{m}_{1}})-\frac{u_{\mathrm{min}}}{u_{\mathrm{max}}}\epsilon(k-w)}$ (119)
conditioned on $\mathcal{E}_{\mathrm{b}}$ for all $w\in\{0,\ldots,k-1\}$, with
$f(\boldsymbol{u}_{\boldsymbol{m}_{1}})=\frac{\|\boldsymbol{u}_{\boldsymbol{m}_{1}}\|_{1}}{u_{\mathrm{max}}}+\frac{u_{\mathrm{max}}\|\boldsymbol{u}_{\boldsymbol{m}_{1}}\|_{1}-\|\boldsymbol{u}_{\boldsymbol{m}_{1}}\|_{2}^{2}}{2\sigma^{2}}.$ (120)
We evaluate $f(\boldsymbol{u}_{\boldsymbol{m}_{1}})$. From $u_{\mathrm{max}}\|\boldsymbol{u}_{\boldsymbol{m}_{1}}\|_{1}-\|\boldsymbol{u}_{\boldsymbol{m}_{1}}\|_{2}^{2}\geq 0$ we use the assumption $2\sigma^{2}>u_{\mathrm{max}}^{2}$ to have
$f(\boldsymbol{u}_{\boldsymbol{m}_{1}})<\frac{2}{u_{\mathrm{max}}}\|\boldsymbol{u}_{\boldsymbol{m}_{1}}\|_{1}-\frac{\|\boldsymbol{u}_{\boldsymbol{m}_{1}}\|_{2}^{2}}{u_{\mathrm{max}}^{2}}.$ (121)
Completing the square with respect to $|\boldsymbol{u}_{\boldsymbol{m}_{1}}|$ yields
$f(\boldsymbol{u}_{\boldsymbol{m}_{1}})<-\frac{\||\boldsymbol{u}_{\boldsymbol{m}_{1}}|-u_{\mathrm{max}}\boldsymbol{1}_{k-w}\|_{2}^{2}}{u_{\mathrm{max}}^{2}}+k-w\leq k-w,$ (122)
where the upper bound is attained at $|\boldsymbol{u}_{\boldsymbol{m}_{1}}|=u_{\mathrm{max}}\boldsymbol{1}$. Applying this upper bound to (119), we arrive at
$G_{k-w}^{k}(\boldsymbol{\omega}_{1})<|\mathcal{M}|^{k-w}\binom{k}{w}\left(N/k\right)^{(1-u_{\mathrm{min}}\epsilon/u_{\mathrm{max}})(k-w)}$ (123)
conditioned on $\mathcal{E}_{\mathrm{b}}$.
We are ready to evaluate $g_{1}$ in (107). Using the upper bound (123) of $G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ conditioned on $\mathcal{E}_{\mathrm{b}}$, we arrive at
$g_{1}$ $<\mathbb{P}(\mathcal{E}_{\mathrm{b}})\log\left\{\sum_{w=0}^{k}\binom{k}{w}\left[\frac{(N/k)^{-u_{\mathrm{min}}\epsilon/u_{\mathrm{max}}}}{1-k/N}\right]^{k-w}\right\}$
$\leq k\log\left\{1+\frac{(N/k)^{-u_{\mathrm{min}}\epsilon/u_{\mathrm{max}}}}{1-k/N}\right\},$ (124)
where the last equality follows from the binomial theorem and $\mathbb{P}(\mathcal{E}_{\mathrm{b}})\leq 1$. The upper bound (124) implies that $g_{1}$ converges to zero in the sublinear sparsity limit for some $\epsilon$ satisfying
$\epsilon\in\left(\frac{\gamma u_{\mathrm{max}}}{(1-\gamma)u_{\mathrm{min}}},\;1-\frac{u_{\mathrm{max}}^{2}}{2\sigma^{2}}\right).$ (125)
Note that $\epsilon$ in (125) exists for
$\gamma<\frac{C_{0}}{C_{0}+\sigma^{2}/(u_{\mathrm{max}}u_{\mathrm{min}})},$ (126)
with $C_{0}$ given in (26).
D-D Non-Typical Noise Samples
For $g_{2}$ in (108), the non-typical event $\mathcal{E}_{\mathrm{b}}^{\mathrm{c}}$ can be represented as $\mathcal{E}_{\mathrm{b}}^{\mathrm{c}}=\cup_{i=1}^{k}(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}$. We use the upper bound $1(\mathcal{E}_{\mathrm{b}}^{\mathrm{c}})\leq\sum_{i=1}^{k}1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})$ to obtain $g_{2}=\sum_{i=1}^{k}g_{2,i}$ with
$g_{2,i}$ $=\mathbb{E}\left[1\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\right)\right$.
$\left.\cdot\log\left\{\sum_{w=0}^{k}[|\mathcal{M}|(N/k-1)]^{w-k}G_{k-w}^{k}(\boldsymbol{\omega}_{1})\right\}\right].$ (127)
It is difficult to evaluate the expectation in (127) with respect to $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$. Furthermore, using Jensen’s inequality results in a loose upper bound on $g_{2,i}$. To circumvent this difficulty, we prove the following upper bound on $G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ in (103) conditioned on $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}$:
Lemma 5
Define
$\mathcal{I}=\{j\in\{1,\ldots,k\}:h(\omega_{j})\leq 1\},$ (128)
with $h(\omega_{j})=h(u_{\mathrm{max}},\omega_{j})$ in (102). Then, we have
$G_{k-w}^{k}(\boldsymbol{\omega}_{1})\leq H_{w}(|\mathcal{I}|)G_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})$ (129)
for all $w\in\{0,\ldots,k-1\}$. In (129), $G_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})\geq 1$ is given in (103). Furthermore, $H_{w}(j)$ is defined as $H_{0}(j)=1$ and
$H_{w}(j)=\sum_{w^{\prime}=\max\{j-w,0\}}^{\min\{j,k-w\}}\binom{j}{w^{\prime}}\lceil L_{k-w-w^{\prime},k-j}^{k-j}\rceil$ (130)
for $w\in\{1,\ldots,k-1\}$, with
$L_{w,w^{\prime}}^{k}=\prod_{i=0}^{w^{\prime}-w-1}\frac{w^{\prime}-i}{k-w-i}.$ (131)
Proof:
See Appendix E. ∎
Lemma 5 allows us to replace intractable $G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ in (127) with $G_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})$ in (103). As shown shortly, $G_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})$ can be bounded from above by a single exponential function of $\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}}$. Thus, an upper bound on $g_{2,i}$ can be evaluated analytically.
We evaluate $g_{2,i}$ in (127). For $w=k$ we have $G_{0}^{k}(\boldsymbol{\omega}_{1})=1\leq G_{k-|\mathcal{I}}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})$ and $H_{k}(j)=\lceil L_{0,k-j}^{k-j}\rceil=1$. For $w\in\{0,\ldots,k-1\}$ we use Lemma 5 to have
$g_{2,i}$ $<\mathbb{E}\left[1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})\log\left\{|\mathcal{M}|^{-(k-|\mathcal{I}|)}G_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})\right\}\right]$
$+\mathbb{E}\left[1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})H(|\mathcal{I}|)\right],$ (132)
with
$H(j)=\log\left\{\sum_{w=0}^{k}|\mathcal{M}|^{w-j}(N/k-1)^{w-k}H_{w}(j)\right\}.$ (133)
The first and second terms on the upper bound (132) are represented as $g_{2,i}^{(1)}$ and $g_{2,i}^{(2)}$, respectively. We use the upper bound (132) to obtain $g_{2}=\sum_{i=1}^{k}g_{2,i}\leq g_{2}^{(1)}+g_{2}^{(2)}$, with $g_{2}^{(l)}=\sum_{i=1}^{k}g_{2,i}^{(l)}$ for $l\in\{1,2\}$. We need to prove that both $g_{2}^{(1)}$ and $g_{2}^{(2)}$ converge to zero in the sublinear sparsity limit.
To evaluate them, we investigate statistical properties of $\mathcal{I}$ in (128). Since $\{\omega_{j}\}$ are i.i.d. random variables, $\{h(\omega_{j})\}$ are also i.i.d. Thus, the distribution of $\mathcal{I}$ depends on $\mathcal{I}$ only through the cardinality $|\mathcal{I}|$. Furthermore, $|\mathcal{I}|\in\{0,\ldots,k\}$ follows the binomial distribution $B(k,p)$ with $p=\mathbb{P}(h(\omega_{1})\leq 1)$: $\mathbb{P}(|\mathcal{I}|=j)=p_{j}^{(k)}$ holds with
$p_{j}^{(k)}=\binom{k}{j}p^{j}(1-p)^{k-j}.$ (134)
In evaluating $g_{2}^{(1)}$ and $g_{2}^{(2)}$, we use two technical results.
Lemma 6
For $\omega\sim\mathcal{N}(0,\sigma_{N/k}^{2})$, we have $\mathbb{P}(h(\omega)\leq 1)\leq(N/k)^{-C_{1,2}}$, with $C_{1,2}$ given in (27).
Proof:
For $\omega\sim\mathcal{N}(0,\sigma_{N/k}^{2})$, we use the definition of $h(\omega)=h(u_{\mathrm{max}},\omega)$ in (102) to find that $h(\omega)\leq 1$ implies
$\frac{|u_{m}|u_{\mathrm{max}}}{\sigma_{N/k}^{2}}-\frac{u_{m}^{2}}{2\sigma_{N/k}^{2}}+\frac{u_{m}\omega}{\sigma_{N/k}^{2}}\leq 0$ (135)
for all $m\in\mathcal{M}$. In particular, selecting $u_{m}=u_{\mathrm{min}}$—without loss of generality $u_{\mathrm{min}}\in\mathcal{U}$ has been assumed—yields $\omega\leq-(u_{\mathrm{max}}-u_{\mathrm{min}}/2)$. We use $\omega\sim\mathcal{N}(0,\sigma_{N/k}^{2})$ to have
$\mathbb{P}(h(\omega)\leq 1)\leq\mathbb{P}\left(\omega\leq-\left(u_{\mathrm{max}}-\frac{u_{\mathrm{min}}}{2}\right)\right)$
$=Q\left(\frac{u_{\mathrm{max}}-u_{\mathrm{min}}/2}{\sigma_{N/k}}\right)\leq(N/k)^{-\frac{(u_{\mathrm{max}}-u_{\mathrm{min}}/2)^{2}}{2\sigma^{2}}},$ (136)
where the last inequality follows from the Chernoff bound for $Q(x)$. Thus, Lemma 6 holds. ∎
Lemma 6 implies that $(1+p)^{k}$ converges to zero in the sublinear sparsity limit for $\gamma<C_{1,2}/(C_{1,2}+1)$.
Proposition 5
For all $x\in\mathbb{R}$, we have
$S(x)=\sum_{j=0}^{k}\binom{k}{j}jx^{j}=kx(1+x)^{k}.$ (137)
Proof:
Define $f(x)=\sum_{j=0}^{k}\binom{k}{j}x^{j}$. Differentiating $f(x)$ yields $f^{\prime}(x)=\sum_{j=1}^{k}\binom{k}{j}jx^{j-1}=x^{-1}S(x)$ or equivalently $S(x)=xf^{\prime}(x)$. Since the binomial theorem implies $f(x)=(1+x)^{k}$, we have $f^{\prime}(x)=k(1+x)^{k-1}$. Thus, $S(x)=xf^{\prime}(x)$ reduces to (137). ∎
We prove the convergence of $g_{2}^{(2)}$ to zero in Appendix D-E while $g_{2}^{(1)}$ is evaluated in Appendix D-F.
D-E Second Term in (132)
For the second term $g_{2,i}^{(2)}$ on the upper bound (132), we evaluate $g_{2}^{(2)}=\sum_{i=1}^{k}g_{2,i}^{(2)}$. We use the representation $g_{2,i}^{(2)}=\sum_{j=0}^{k}g_{2,i,j}^{(2)}$ with
$g_{2,i,j}^{(2)}=\mathbb{P}\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\right)H(j).$ (138)
We first evaluate $H(j)$ in (133). For $w\in\{1,\ldots,k-1\}$, we use $\max\{j-w,0\}\geq 0$ and $\min\{j,k-w\}\leq j$ to upper-bound $H_{w}(j)$ in (130) as
$H_{w}(j)\leq\sum_{w^{\prime}=0}^{j}\binom{j}{w^{\prime}}\lceil L_{k-w-w^{\prime},k-j}^{k-j}\rceil,$ (139)
with
$L_{k-w-w^{\prime},k-j}^{k-j}=\prod_{i=0}^{w+w^{\prime}-j-1}\frac{k-j-i}{w+w^{\prime}-j-i}<k^{w+w^{\prime}-j}.$ (140)
The binomial theorem implies
$H_{w}(j)<k^{w-j}(k+1)^{j}.$ (141)
In particular, (141) is also correct for $w=0$ and $w=k$, because of $H_{0}(j)=H_{k}(j)=1$. We substitute the upper bound (141) into $H(j)$ in (133) to arrive at
$H(j)$ $<\log\left\{\frac{(|\mathcal{M}|k)^{-j}(k+1)^{j}}{(N/k-1)^{k}}\frac{\{|\mathcal{M}|(N-k)\}^{k+1}-1}{|\mathcal{M}|(N-k)-1}\right\}$
$<j\log\left(\frac{k+1}{|\mathcal{M}|k}\right)+\log\frac{(|\mathcal{M}|k)^{k+1}(N/k-1)}{|\mathcal{M}|k(N/k-1)-1}.$ (142)
We next focus on $\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\})$, in which the probability distribution is induced from $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$ through $\mathcal{I}$ in (128) and $\mathcal{E}_{i}^{k-1}=\{\Omega_{i}^{k-1}:|\Omega_{i}^{k-1}|\leq d_{i}^{k-1}\}$ with $\Omega_{i}^{k-1}=(\boldsymbol{x}_{i}^{k-1})^{\mathrm{T}}\boldsymbol{\omega}_{1}/\sigma_{N/k}^{2}$. Decomposing the event $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ yields
$\mathbb{P}\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\right)$
$=\mathbb{P}\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\in\mathcal{I}\}\right)$
$+\mathbb{P}\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\notin\mathcal{I}\}\right).$ (143)
We evaluate the first term in (143). Since $h(\omega_{i^{\prime}})\leq 1$ holds for all $i^{\prime}\in\mathcal{I}\setminus\{i\}$, as well as $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{h(\omega_{i})\leq 1\}$, we have $\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=0\}\cap\{i\in\mathcal{I}\})=0$, $\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=1\}\cap\{i\in\mathcal{I}\})=\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{h(\omega_{i})\leq 1\})\leq\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})$, and
$\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\in\mathcal{I}\})$
$\leq p_{j-1}^{(k-1)}\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})\leq\binom{k-1}{j-1}p^{j-1}\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})$ (144)
for all $j\geq 2$, with $p_{j}^{(k)}$ defined in (134) and $p=\mathbb{P}(h(\omega_{1})\leq 1)$. In particular, this upper bound is also correct for $j=1$. For the second term in (143), similarly, we obtain $\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=k\}\cap\{i\notin\mathcal{I}\})=0$ and
$\mathbb{P}\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\notin\mathcal{I}\}\right)$
$\leq p_{j}^{(k-1)}\mathbb{P}\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\right)\leq\binom{k-1}{j}p^{j}\mathbb{P}\left((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\right).$ (145)
for all $j<k$.
We are ready to evaluate $g_{2,i}^{(2)}=\sum_{j=0}^{k}g_{2,i,j}^{(2)}$. Substituting the upper bounds (142), (144), and (145) into $g_{2,i,j}^{(2)}$ in (138), from the binomial theorem we obtain
$\frac{g_{2,i}^{(2)}}{\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})}<\left\{\sum_{j=1}^{k}\binom{k-1}{j-1}jp^{j-1}+\sum_{j=0}^{k-1}\binom{k-1}{j}jp^{j}\right\}$
$\cdot\log\left(\frac{k+1}{|\mathcal{M}|k}\right)+2(1+p)^{k-1}\log\frac{(|\mathcal{M}|k)^{k+1}(N/k-1)}{|\mathcal{M}|k(N/k-1)-1}.$ (146)
Applying Proposition 5 to the upper bound (146) and using Lemmas 4 and 6, we find that $g_{2}^{(2)}=\sum_{i=1}^{k}g_{2,i}^{(2)}$ reduces to
$g_{2}^{(2)}<\frac{2ke^{C_{0}\epsilon}}{(N/k)^{C_{1,1}}}\left\{\left[2(k-1)(N/k)^{-C_{1,2}}+1\right]\log\left(\frac{k+1}{|\mathcal{M}|k}\right)\right$.
$\left.+2\log\frac{(|\mathcal{M}|k)^{k+1}(N/k-1)}{|\mathcal{M}|k(N/k-1)-1}\right\}\left\{1+(N/k)^{-C_{1,2}}\right\}^{k-1}.$ (147)
This upper bound implies that $g_{2}^{(2)}$ converges to zero in the sublinear sparsity limit for $\epsilon$ satisfying (125) and
$\gamma$ $<\min\left\{\frac{C_{1,1}+2C_{1,2}}{C_{1,1}+2C_{1,2}+3},\;\frac{C_{1,1}+C_{1,2}}{C_{1,1}+C_{1,2}+3}\right\}$
$=\frac{C_{1,1}+C_{1,2}}{C_{1,1}+C_{1,2}+3}.$ (148)
D-F First Term in (132)
For the first term $g_{2,i}^{(1)}$ on the upper bound (132), we evaluate $g_{2}^{(1)}=\sum_{i=1}^{k}g_{2,i}^{(1)}$. We use the representation $g_{2,i}^{(1)}=\sum_{j=0}^{k}g_{2,i,j}^{(1)}$ with
$g_{2,i,j}^{(1)}=\mathbb{E}\left[1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\})\log\left\{|\mathcal{M}|^{-(k-j)}G_{k-j}^{k-j}\right\}\right],$ (149)
where the probability distribution is induced from $\boldsymbol{\omega}_{1}\sim\mathcal{N}(\boldsymbol{0},\sigma_{N/k}^{2}\boldsymbol{I}_{k})$ through $\mathcal{I}$ in (128) and $\mathcal{E}_{i}^{k-1}=\{\Omega_{i}^{k-1}:|\Omega_{i}^{k-1}|\leq d_{i}^{k-1}\}$ with $\Omega_{i}^{k-1}=(\boldsymbol{x}_{i}^{k-1})^{\mathrm{T}}\boldsymbol{\omega}_{1}/\sigma_{N/k}^{2}$.
We decompose $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ into two disjoint events $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\in\mathcal{I}\}$ and $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\notin\mathcal{I}\}$. Representing $g_{2,i,j}^{(1)}$ with $(\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}$ replaced by the former and latter events as $g_{2,i,j}^{(1,1)}$ and $g_{2,i,j}^{(1,2)}$, respectively, we have $g_{2,i,j}^{(1)}=g_{2,i,j}^{(1,1)}+g_{2,i,j}^{(1,2)}$. The goal is evaluation of $g_{2}^{(1)}=g_{2}^{(1,1)}+g_{2}^{(1,2)}$ with $g_{2}^{(1,l)}=\sum_{i=1}^{k}g_{2,i}^{(1,l)}$ and $g_{2,i}^{(1,l)}=\sum_{j=0}^{k}g_{2,i,j}^{(1,l)}$ for $l\in\{1,2\}$.
For $g_{2,i,j}^{(1,1)}$, we use the definition of $G_{k-j}^{k-j}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})$ in (103) and Jensen’s inequality to obtain
$g_{2,i,j}^{(1,1)}$ $\leq\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\in\mathcal{I}\})$
$\cdot\log\left\{\frac{1}{|\mathcal{M}|^{k-j}}\sum_{\boldsymbol{m}\in\mathcal{M}^{k-j}}e^{u_{\mathrm{max}}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{1}/\sigma_{N/k}^{2}}\right\}.$ (150)
From $\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=0\}\cap\{i\in\mathcal{I}\})=0$, we have $g_{2,i,0}^{(1,1)}=0$. For all $j\in\{1,\ldots,k\}$, we use the upper bound (144) and $\|\boldsymbol{u}_{\boldsymbol{m}}\|_{1}\leq u_{\mathrm{max}}(k-j)$ to obtain
$g_{2,i,j}^{(1,1)}\leq\frac{u_{\mathrm{max}}^{2}}{\sigma_{N/k}^{2}}(k-j)\binom{k-1}{j-1}p^{j-1}\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}).$ (151)
Utilizing Proposition 5 and the binomial theorem, we find that $g_{2,i}^{(1,1)}=\sum_{j=1}^{k}g_{2,i,j}^{(1,1)}$ is bounded from above by
$g_{2,i}^{(1,1)}$ $\leq\frac{u_{\mathrm{max}}^{2}}{\sigma_{N/k}^{2}}\left\{k\sum_{j=1}^{k}\binom{k-1}{j-1}p^{j-1}-\sum_{j=1}^{k}\binom{k-1}{j-1}jp^{j-1}\right\}$
$\cdot\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})$
$=\frac{u_{\mathrm{max}}^{2}(k-1)}{\sigma_{N/k}^{2}}\left(1-p\right)\left(1+p\right)^{k-1}\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}).$ (152)
Applying Lemmas 4 and 6 to this upper bound, we have the following upper bound on $g_{2}^{(1,1)}=\sum_{i=1}^{k}g_{2,i}^{(1,1)}$:
$g_{2}^{(1,1)}<\frac{2u_{\mathrm{max}}^{2}k^{2}}{\sigma_{N/k}^{2}}\left\{1+(N/k)^{-C_{1,2}}\right\}^{k-1}\frac{e^{C_{0}\epsilon}}{(N/k)^{C_{1,1}}}.$ (153)
This upper bound implies that $g_{2}^{(1,1)}$ converges to zero in the sublinear sparsity limit for $\epsilon$ satisfying (125) and
$\gamma<\frac{C_{1,1}+C_{1,2}}{C_{1,1}+C_{1,2}+3}.$ (154)
For $g_{2,i,j}^{(1,2)}$, similarly, we use the definition of $G_{k-j}^{k-j}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})$ in (103) and Jensen’s inequality with respect to $\{\omega_{i^{\prime}}:i^{\prime}\in\mathcal{I}^{\mathrm{c}}\setminus\{i\}\}$ to obtain
$g_{2,i,j}^{(1,2)}$ $\leq\mathbb{E}\left[1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}}\cap\{|\mathcal{I}|=j\}\cap\{i\notin\mathcal{I}\})\right$.
$\left.\cdot\log\left\{\frac{1}{|\mathcal{M}|^{k-j}}\sum_{\boldsymbol{m}\in\mathcal{M}^{k-j}}F_{i,\boldsymbol{m}}(\omega_{i})\right\}\right],$ (155)
with
$F_{i,\boldsymbol{m}}(\omega)=\exp\left(\frac{u_{\mathrm{max}}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{1}-2^{-1}u_{m_{i}}^{2}+u_{m_{i}}\omega}{\sigma_{N/k}^{2}}\right).$ (156)
By definition, $g_{2,i,k}^{(1,2)}=0$ and
$g_{2,i,j}^{(1,2)}\leq p_{j}^{(k-1)}\mathbb{E}\left[1((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})\log\left\{\sum_{\boldsymbol{m}\in\mathcal{M}^{k-j}}\frac{F_{i,\boldsymbol{m}}(\omega_{i})}{|\mathcal{M}|^{k-j}}\right\}\right]$ (157)
hold for all $j<k$, with $p_{j}^{(k)}$ defined in (134). Using the upper bound $F_{i,\boldsymbol{m}}(\omega)<e^{\{u_{\mathrm{max}}^{2}(k-j)+u_{m_{i}}\omega\}/\sigma_{N/k}^{2}}$ and the definition $\Omega_{i}^{k-1}=u_{*}\omega_{i}/\sigma_{N/k}^{2}$ with $|u_{*}|=u_{\mathrm{max}}$, as well as $d_{i}^{k-1}=d_{*}$, we obtain
$\frac{g_{2,i,j}^{(1,2)}}{p_{j}^{(k-1)}}-\frac{u_{\mathrm{max}}^{2}(k-j)}{\sigma_{N/k}^{2}}\mathbb{P}((\mathcal{E}_{i}^{k-1})^{\mathrm{c}})$
$<\int_{|\Omega_{i}^{k-1}|>d_{*}}\log\left(\sum_{\boldsymbol{m}\in\mathcal{M}^{k-j}}\frac{e^{\frac{u_{m_{i}}\Omega_{i}^{k-1}}{u_{*}}}}{|\mathcal{M}|^{k-j}}\right)p(\Omega_{i}^{k-1})d\Omega_{i}^{k-1}$
$\leq 2\int_{d_{*}}^{\infty}\Omega_{i}^{k-1}p(\Omega_{i}^{k-1})d\Omega_{i}^{k-1}=\frac{u_{\mathrm{max}}}{\sigma_{N/k}}\sqrt{\frac{2}{\pi}}e^{-\frac{\sigma_{N/k}^{2}d_{*}^{2}}{2u_{\mathrm{max}}^{2}}}$
$\leq\frac{u_{\mathrm{max}}}{\sigma_{N/k}}\sqrt{\frac{2}{\pi}}\frac{e^{C_{0}\epsilon}}{(N/k)^{C_{1,1}}}.$ (158)
In the derivation of the second inequality, we have used $|u_{m_{i}}\Omega_{i}^{k-1}/u_{*}|\leq|\Omega_{i}^{k-1}|$ while the equality is obtained from $\Omega_{i}^{k-1}\sim\mathcal{N}(0,u_{\mathrm{max}}^{2}/\sigma_{N/k}^{2})$. The last inequality follows from the lower bound (116) on $d_{*}$. Thus, we use Lemma 4 to arrive at
$g_{2,i,j}^{(1,2)}$ $<p_{j}^{(k-1)}\left\{\frac{2u_{\mathrm{max}}^{2}(k-j)}{\sigma_{N/k}^{2}}+\frac{u_{\mathrm{max}}}{\sigma_{N/k}}\sqrt{\frac{2}{\pi}}\right\}\frac{e^{C_{0}\epsilon}}{(N/k)^{C_{1,1}}}.$ (159)
We are ready to evaluate $g_{2}^{(1,2)}=\sum_{i=1}^{k}g_{2,i}^{(1,2)}$ with $g_{2,i}^{(1,2)}=\sum_{j=0}^{k-1}g_{2,i,j}^{(1,2)}$. Utilizing $p_{j}^{(k-1)}\leq\binom{k-1}{j}p^{j}$ in (134), the binomial theorem, and Proposition 5, we have
$\frac{g_{2}^{(1,2)}}{ke^{C_{0}\epsilon}}<\left\{\frac{2u_{\mathrm{max}}^{2}[k-(k-1)p]}{\sigma_{N/k}^{2}}+\frac{u_{\mathrm{max}}}{\sigma_{N/k}}\sqrt{\frac{2}{\pi}}\right\}\frac{(1+p)^{k-1}}{(N/k)^{C_{1,1}}}.$ (160)
We use Lemma 6 to arrive at
$g_{2}^{(1,2)}$ $<\left(\frac{2u_{\mathrm{max}}^{2}k^{2}}{\sigma_{N/k}^{2}}+\frac{u_{\mathrm{max}}k}{\sigma_{N/k}}\sqrt{\frac{2}{\pi}}\right)$
$\cdot\left\{1+(N/k)^{-C_{1,2}}\right\}^{k-1}\frac{e^{C_{0}\epsilon}}{(N/k)^{C_{1,1}}}.$ (161)
This upper bound implies that $g_{2}^{(1,2)}$ converges to zero in the sublinear sparsity limit for $\epsilon$ satisfying (125) and
$\gamma<\frac{C_{1,1}+C_{1,2}}{C_{1,1}+C_{1,2}+3}.$ (162)
Thus, Lemma 2 holds.
Appendix E Proof of Lemma 5
For $w=0$, we represent $G_{k}^{k}(\boldsymbol{\omega}_{1})$ in (103) as
$G_{k}^{k}(\boldsymbol{\omega}_{1})=\prod_{j=1}^{k}h(\omega_{j}),$ (163)
with $h(\omega)=h(u_{\mathrm{max}},\omega)$ in (102). Thus, we use $\mathcal{I}=\{j\in\{1,\ldots,k\}:h(\omega_{j})\leq 1\}$ to find $G_{k}^{k}(\boldsymbol{\omega}_{1})\leq G_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}})$. In particular, from $H_{0}(j)=1$ we obtain the upper bound (129) for $w=0$.
For $w\in\{1,\ldots,k-1\}$, we consider the following decomposition of $\mathcal{X}_{k-w}^{k}(\mathcal{U})$:
$\mathcal{X}_{k-w}^{k}(\mathcal{U})=\bigcup_{w^{\prime}=\max\{|\mathcal{I}|-w,0\}}^{\min\{|\mathcal{I}|,k-w\}}\mathcal{X}_{w^{\prime}}^{|\mathcal{I}|}(\mathcal{U})\times\mathcal{X}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\mathcal{U}),$ (164)
where $\mathcal{X}_{w^{\prime}}^{|\mathcal{I}|}(\mathcal{U})$ corresponds to vectors in positions $\mathcal{I}$ while $\mathcal{X}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\mathcal{U})$ is associated with $\mathcal{I}^{\mathrm{c}}$. We use this decomposition to represent $G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ in (103) as
$G_{k-w}^{k}(\boldsymbol{\omega}_{1})$ $=\sum_{w^{\prime}=\max\{|\mathcal{I}|-w,0\}}^{\min\{|\mathcal{I}|,k-w\}}\tilde{G}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,2})$
$\cdot\sum_{\boldsymbol{x}_{1,1}^{\prime}\in\mathcal{X}_{w^{\prime}}^{|\mathcal{I}|}(\{1\})}\prod_{j\in\mathcal{S}_{\boldsymbol{x}_{1,1}^{\prime}}}h([\boldsymbol{\omega}_{1,1}]_{j}),$ (165)
with
$\tilde{G}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,2})=\sum_{\boldsymbol{x}_{1,2}^{\prime}\in\mathcal{X}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\{1\})}\prod_{j\in\mathcal{S}_{\boldsymbol{x}_{1,2}^{\prime}}}h([\boldsymbol{\omega}_{1,2}]_{j}),$ (166)
where $\boldsymbol{\omega}_{1}\in\mathbb{R}^{k}$ satisfies $\boldsymbol{\omega}_{1,\mathcal{I}}=\boldsymbol{\omega}_{1,1}$ and $\boldsymbol{\omega}_{1,\mathcal{I}^{\mathrm{c}}}=\boldsymbol{\omega}_{1,2}$. From the definition of $\mathcal{I}$, as well as $h(\omega_{j})=h(u_{\mathrm{max}},\omega_{j})$ in (102), we obtain
$G_{k-w}^{k}(\boldsymbol{\omega}_{1})\leq\sum_{w^{\prime}=\max\{|\mathcal{I}|-w,0\}}^{\min\{|\mathcal{I}|,k-w\}}\binom{|\mathcal{I}|}{w^{\prime}}\tilde{G}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,2}).$ (167)
In particular, we use the definition of $G_{k}^{k}$ in (103) to find $\tilde{G}_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,2})=G_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,2})\geq 1$.
To upper-bound $\tilde{G}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,2})$, we represent (166) with $h(\omega)=h(u_{\mathrm{max}},\omega)$ in (102) as
$\tilde{G}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\boldsymbol{\omega}_{1,2})=\sum_{\boldsymbol{x}_{1,2}^{\prime}\in\mathcal{X}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}(\mathcal{U})}\prod_{i=1}^{k-|\mathcal{I}|}f_{i}([\boldsymbol{x}_{1,2}^{\prime}]_{i}),$ (168)
with
$f_{i}(x)=\exp\left(\frac{u_{\mathrm{max}}|x|}{\sigma_{N/k}^{2}}-\frac{|x|^{2}}{2\sigma_{N/k}^{2}}+\frac{x[\boldsymbol{\omega}_{1,2}]_{i}}{\sigma_{N/k}^{2}}\right).$ (169)
Lemma 7
For $i\in\{1,\ldots,k\}$, suppose that $f_{i}:\mathbb{R}\to[0,\infty)$ satisfies $f_{i}(0)\leq\sum_{m\in\mathcal{M}}f_{i}(u_{m})$. Then, we have
$\sum_{\boldsymbol{x}\in\mathcal{X}_{w}^{k}(\mathcal{U})}\prod_{i=1}^{k}f_{i}(x_{i})\leq\lceil L_{w,w^{\prime}}^{k}\rceil\sum_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{w^{\prime}}^{k}(\mathcal{U})}\prod_{i=1}^{k}f_{i}(x_{i}^{\prime})$ (170)
for $w<w^{\prime}$, with $L_{w,w^{\prime}}^{k}$ defined in (131).
Proof:
See Appendix F. ∎
From the condition $h(\omega_{i})>1$ for $i\notin\mathcal{I}$, as well as the definition of $h(\omega_{i})=h(u_{\mathrm{max}},\omega_{i})$ in (102), we find $f_{i}(0)=1$ and $\sum_{m\in\mathcal{M}}f_{i}(u_{m})=h(\omega_{i})>1$ for $f_{i}$ in (169). Thus, we can use Lemma 7 to obtain $\tilde{G}_{k-w-w^{\prime}}^{k-|\mathcal{I}|}\leq\lceil L_{k-w-w^{\prime},k-|\mathcal{I}|}^{k-|\mathcal{I}|}\rceil\tilde{G}_{k-|\mathcal{I}|}^{k-|\mathcal{I}|}$. Combining these results, we arrive at the upper bound (129).
Appendix F Proof of Lemma 7
We introduce a graphical relationship between $\mathcal{X}_{w}^{k}(\mathcal{U})$ and $\mathcal{X}_{w^{\prime}}^{k}(\mathcal{U})$ for $w<w^{\prime}$. Let $\mathcal{V}_{w}^{k}=\mathcal{X}_{w}^{k}(\{1\})$ denote the set of all $k$-dimensional, $w$-sparse, and binary vectors. We consider an undirected bipartite graph $G(\mathcal{V}_{w}^{k}\cup\mathcal{V}_{w^{\prime}}^{k},\mathcal{E})$, in which there is an edge $(v,v^{\prime})\in\mathcal{E}$ between $v\in\mathcal{V}_{w}^{k}$ and $v^{\prime}\in\mathcal{V}_{w^{\prime}}^{k}$ only when the support $\mathcal{S}_{v}$ of the $w$-sparse binary vector $v\in\mathcal{X}_{w}^{k}(\{1\})$ is a subset of the support $\mathcal{S}_{v^{\prime}}$ of $v^{\prime}\in\mathcal{X}_{w^{\prime}}^{k}(\{1\})$.
The significance of this edge definition can be understood as follows: For an edge $(v,v^{\prime})\in\mathcal{E}=\mathcal{V}_{w}^{k}\times\mathcal{V}_{w^{\prime}}^{k}$, suppose that $\boldsymbol{x}\in\mathcal{X}_{w}^{k}(\mathcal{U})$ and $\boldsymbol{x}^{\prime}\in\mathcal{X}_{w^{\prime}}^{k}(\mathcal{U})$ satisfy $\mathcal{S}_{\boldsymbol{x}}=\mathcal{S}_{v}$, $\mathcal{S}_{\boldsymbol{x}^{\prime}}=\mathcal{S}_{v^{\prime}}$, and $x_{i}=x_{i}^{\prime}$ for all $i\in\mathcal{S}_{\boldsymbol{x}}$. By definition, each term in (170) satisfies
$\prod_{i=1}^{k}f_{i}(x_{i})\leq\sum_{i^{\prime}\in\mathcal{S}_{\boldsymbol{x}^{\prime}}\setminus\mathcal{S}_{\boldsymbol{x}}}\sum_{x_{i^{\prime}}^{\prime}\in\mathcal{U}}\prod_{i=1}^{k}f_{i}(x_{i}^{\prime}),$ (171)
because of the assumption $f_{i}(0)\leq\sum_{m\in\mathcal{M}}f_{i}(u_{m})$ for all $i$. Thus, we can upper-bound the left-hand side (LHS) in (170) by replacing $\boldsymbol{x}$ with $\boldsymbol{x}^{\prime}$, if there is a one-to-one correspondence between $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$. To prove the existence of a one-to-one correspondence, we need the notion of a perfect matching in graph theory.
Definition 2
For $w<w^{\prime}\leq\lceil k/2\rceil$, a subset $\mathcal{M}\subset\mathcal{E}$ of edges in the undirected bipartite graph $G(\mathcal{V}_{w}^{k}\cup\mathcal{V}_{w^{\prime}}^{k},\mathcal{E})$ is called a perfect matching with respect to $\mathcal{V}_{w}^{k}$ if $\mathcal{M}$ has $|\mathcal{V}_{w}^{k}|$ elements and if different edges in $\mathcal{M}$ are not adjacent to each other: Different edges $(v_{1},v_{1}^{\prime}),(v_{2},v_{2}^{\prime})\in\mathcal{M}$ satisfy $v_{1}\neq v_{2}$ and $v_{2}\neq v_{2}^{\prime}$.
When the graph $G(\mathcal{V}_{w}^{k}\cup\mathcal{V}_{w^{\prime}}^{k},\mathcal{E})$ has a perfect matching with respect to $\mathcal{V}_{w}^{k}$, there is a one-to-one correspondence between $\boldsymbol{x}\in\mathcal{X}_{w}^{k}(\mathcal{U})$ and $\boldsymbol{x}^{\prime}\in\mathcal{X}_{w^{\prime}}^{k}(\mathcal{U})$. Thus, we have the upper bound (170) with $L_{w,w^{\prime}}^{k}=1$.
To prove the existence of a perfect matching, we utilize a general result for a regular bipartite graph $G(\mathcal{V}_{1}\cup\mathcal{V}_{2},\mathcal{E})$. Let $\mathcal{N}_{1}(v)\subset\mathcal{V}_{2}$ denote the neighborhood of $v\in\mathcal{V}_{1}$, i.e. the set of all $v^{\prime}\in\mathcal{V}_{2}$ adjacent to $v$. Similarly, the neighborhood of $v^{\prime}\in\mathcal{V}_{2}$ is written as $\mathcal{N}_{2}(v^{\prime})\subset\mathcal{V}_{1}$. The graph is said to be $(d_{1},d_{2})$-regular when $|\mathcal{N}_{1}(v)|=d_{1}$ and $|\mathcal{N}_{2}(v^{\prime})|=d_{2}$ are satisfied for all $v\in\mathcal{V}_{1}$ and $v^{\prime}\in\mathcal{V}_{2}$. Since the number of outgoing edges from $\mathcal{V}_{1}$ is equal to that of incoming edges to $\mathcal{V}_{2}$, we have $d_{1}|\mathcal{V}_{1}|=d_{2}|\mathcal{V}_{2}|$.
Lemma 8
Suppose that $G(\mathcal{V}_{1}\cup\mathcal{V}_{2},\mathcal{E})$ is $(d_{1},d_{2})$-regular. Then, there is a perfect matching with respect to $\mathcal{V}_{1}$ if and only if $|\mathcal{V}_{1}|\leq|\mathcal{V}_{2}|$ holds.
Proof:
Hall’s marriage theorem [41] implies that there is a perfect matching with respect to $\mathcal{V}_{1}$ in $G(\mathcal{V}_{1}\cup\mathcal{V}_{2},\mathcal{E})$ if and only if all subsets $\mathcal{V}\subset\mathcal{V}_{1}$ satisfy the marriage condition $|\mathcal{V}|\leq|\mathcal{N}_{1}(\mathcal{V})|$ with $\mathcal{N}_{1}(\mathcal{V})=\cup_{v\in\mathcal{V}}\mathcal{N}_{1}(v)$. Thus, it is sufficient to confirm that the marriage condition is equivalent to $|\mathcal{V}_{1}|\leq|\mathcal{V}_{2}|$. The marriage condition for $\mathcal{V}=\mathcal{V}_{1}$ implies $|\mathcal{V}_{1}|\leq|\mathcal{V}_{2}|$.
We prove that $|\mathcal{V}_{1}|\leq|\mathcal{V}_{2}|$ implies the marriage condition. For any subset $\mathcal{V}\subset\mathcal{V}_{1}$, there are $d_{1}|\mathcal{V}|$ edges that are outgoing from $\mathcal{V}$ to $\mathcal{N}_{1}(\mathcal{V})$. Since $v^{\prime}\in\mathcal{V}_{2}$ has $d_{2}$ edges, we have the inequality $d_{1}|\mathcal{V}|\leq d_{2}|\mathcal{N}_{1}(\mathcal{V})|$. From the identity $d_{1}|\mathcal{V}_{1}|=d_{2}|\mathcal{V}_{2}|$ and the assumption $|\mathcal{V}_{1}|\leq|\mathcal{V}_{2}|$, we find $d_{1}\geq d_{2}$. Combining these two inequalities, we arrive at the marriage condition $|\mathcal{N}_{1}(\mathcal{V})|\geq(d_{1}/d_{2})|\mathcal{V}|\geq|\mathcal{V}|$. ∎
Corollary 1
If $|\mathcal{V}_{w}^{k}|\leq|\mathcal{V}_{w^{\prime}}^{k}|$ holds, then the graph $G(\mathcal{V}_{w}^{k}\cup\mathcal{V}_{w^{\prime}}^{k},\mathcal{E})$ has a perfect matching with respect to $\mathcal{V}_{w}^{k}$.
Proof:
From the assumption $|\mathcal{V}_{w}^{k}|\leq|\mathcal{V}_{w^{\prime}}^{k}|$ and Lemma 8, it is sufficient to confirm that $G(\mathcal{V}_{w}^{k}\cup\mathcal{V}_{w^{\prime}}^{k},\mathcal{E})$ is regular.
For any $v\in\mathcal{V}_{w}^{k}$, the neighborhood $\mathcal{N}_{w}(v)$ of $v$ consists of $v^{\prime}\in\mathcal{V}_{w^{\prime}}^{k}$ with $\mathcal{S}_{v}\subset\mathcal{S}_{v^{\prime}}$. In other words, $v^{\prime}\in\mathcal{V}_{w^{\prime}}^{k}$ is obtained by flipping $w^{\prime}-w$ zeros in the $k$-dimensional $w$-sparse vector $v$ to $1$. Thus, $|\mathcal{N}_{w}(v)|=\binom{k-w}{w^{\prime}-w}$ holds. Similarly, we have $|\mathcal{N}_{w^{\prime}}(v^{\prime})|=\binom{w^{\prime}}{w^{\prime}-w}$ for the neighborhood $\mathcal{N}_{w^{\prime}}(v^{\prime})$ of $v^{\prime}\in\mathcal{V}_{w^{\prime}}^{k}$. These results imply that $G(\mathcal{V}_{w}^{k}\cup\mathcal{V}_{w^{\prime}}^{k},\mathcal{E})$ is $(d_{1},d_{2})$-regular with $d_{1}=\binom{k-w}{w^{\prime}-w}$ and $d_{2}=\binom{w^{\prime}}{w^{\prime}-w}$. ∎
[…]
Lemma 9
[…]
Proof:
[…]
$\frac{|\tilde{\mathcal{V}}_{w^{\prime}}^{k}|}{|\mathcal{V}_{w}^{k}|}$ $=\lceil L_{w,w^{\prime}}^{k}\rceil\binom{k}{w^{\prime}}\binom{k}{w}^{-1}$
$=\frac{\lceil L_{w,w^{\prime}}^{k}\rceil(k-w)\ldots(k-w^{\prime}+1)}{w^{\prime}\cdots(w+1)}\geq 1,$ (172)
where the inequality follows from the definition of $L_{w,w^{\prime}}^{k}$ in (131). Thus, Lemma 9 holds. ∎
Lemma 9 contains Corollary 1, because of $L_{w,w^{\prime}}^{k}\leq 1$ for $|\mathcal{V}_{w}^{k}|\leq|\mathcal{V}_{w^{\prime}}^{k}|$. Furthermore, the constant $L_{w,w^{\prime}}^{k}$ is tight because $(\lceil L_{w,w^{\prime}}^{k}\rceil-1)|\mathcal{V}_{w^{\prime}}^{k}|<|\mathcal{V}_{w}^{k}|$ holds.
Appendix G Proof of Lemma 3
We know that the KL divergence can be represented with mutual information. To connect the mutual information with the square error, we focus on an identity between them [37]. In particular, we utilize the identity by introducing a perturbation channel,
$y_{N+1}=(\alpha/k)^{1/2}\boldsymbol{a}^{\mathrm{T}}\boldsymbol{x}+\omega_{N+1},\quad\omega_{N+1}\sim\mathcal{N}(0,1).$ (173)
In (173), $\boldsymbol{a}\in\mathbb{R}^{N}$ has independent standard Gaussian elements while the noise $\omega_{N+1}$ is independent of $\{\boldsymbol{x},\boldsymbol{y},\boldsymbol{a}\}$. We recall that $\boldsymbol{x}$ is uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$, while $\boldsymbol{y}$ is defined in the AWGN channel (1). The coefficient $\alpha>0$ has been introduced to control SNR in the perturbation channel. The limit $\alpha\to 0$ is considered to eliminate the influence of the perturbation channel from the final result. The overall channel is represented with $\tilde{\boldsymbol{y}}=[\boldsymbol{y}^{\mathrm{T}},y_{N+1}]^{\mathrm{T}}$.
We first connect $I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})$ with the square error. Consider the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})=\mathbb{E}[\boldsymbol{x}|\boldsymbol{y}]$ of $\boldsymbol{x}$ given $\boldsymbol{y}$. Since $\{\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),y_{N+1}\}$ is a sufficient statistic in estimating $\boldsymbol{x}$ from $\tilde{\boldsymbol{y}}$ conditioned on $\boldsymbol{a}$, we have $I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})=I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),y_{N+1}|\boldsymbol{a})$. Using the chain rule for mutual information [39], we have
$I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})=I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})|\boldsymbol{a})+I(\boldsymbol{x};y_{N+1}|\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),\boldsymbol{a}).$ (174)
For the first term, we obtain
$I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})|\boldsymbol{a})=I(\boldsymbol{x};\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}))=I(\boldsymbol{x};\boldsymbol{y})\leq C_{\mathrm{AWGN}},$ (175)
where $C_{\mathrm{AWGN}}$ denotes the capacity (28) of the AWGN channel (1).
We evaluate the second term in (174). The mutual information $I(\boldsymbol{x};y_{N+1}|\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),\boldsymbol{a})$ is equivalent to the mutual information between $\Delta\boldsymbol{x}=\boldsymbol{x}-\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})$ and $y_{N+1}-(\alpha/k)^{1/2}\boldsymbol{a}^{\mathrm{T}}\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})=(\alpha/k)^{1/2}\boldsymbol{a}^{\mathrm{T}}\Delta\boldsymbol{x}+\omega_{N+1}$ conditioned on $\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})$ and $\boldsymbol{a}$. Replacing $\Delta\boldsymbol{x}$ with capacity-achieving signaling [42]—a zero-mean Gaussian vector with covariance $N^{-1}\mathbb{E}[\|\Delta\boldsymbol{x}\|_{2}^{2}]\boldsymbol{I}_{N}$, we have
$I(\boldsymbol{x};y_{N+1}|\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),\boldsymbol{a})$ $\leq\frac{1}{2}\mathbb{E}\left[\log\left(1+\frac{\alpha\|\boldsymbol{a}\|_{2}^{2}\mathbb{E}[\|\Delta\boldsymbol{x}\|_{2}^{2}]}{kN}\right)\right]$
$\leq\frac{1}{2}\log\left(1+\frac{\alpha}{k}\mathbb{E}[\|\Delta\boldsymbol{x}\|_{2}^{2}]\right),$ (176)
where the last inequality follows from Jensen’s inequality and $\boldsymbol{a}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{N})$. Using the fact that the posterior mean estimator $\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y})$ minimizes the mean-square error, we arrive at
$I(\boldsymbol{x};y_{N+1}|\hat{\boldsymbol{x}}_{\mathrm{opt}}(\boldsymbol{y}),\boldsymbol{a})\leq\frac{1}{2}\log\left\{1+\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}]\right\}$ (177)
for any estimator $\hat{\boldsymbol{x}}(\boldsymbol{y})$ of $\boldsymbol{x}$ based on the observation $\boldsymbol{y}$.
We next connect the mutual information with the KL divergence. For any pdf $q(\boldsymbol{y})$, we use the factorization $p(\tilde{\boldsymbol{y}}|\boldsymbol{x},\boldsymbol{a})=p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})$ to represent the conditional mutual information $I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})$ as
$I(\boldsymbol{x};\tilde{\boldsymbol{y}}|\boldsymbol{a})$ $=\mathbb{E}\left[\log\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})}{q(\boldsymbol{y})}\right]+\mathbb{E}\left[\log\frac{p(\boldsymbol{y})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})}{p(\tilde{\boldsymbol{y}}|\boldsymbol{a})}\right]$
$-D(p\|q),$ (178)
with $p(\boldsymbol{y})=\mathbb{E}_{\boldsymbol{x}}[p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})]$ and $p(\tilde{\boldsymbol{y}}|\boldsymbol{a})=\mathbb{E}_{\boldsymbol{x}}[p(\tilde{\boldsymbol{y}}|\boldsymbol{x},\boldsymbol{a})]$. Combining (174), (175), (177), and (178), we arrive at
$\frac{1}{2}\log\left\{1+\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}\right\}$
$\geq\mathbb{E}\left[\log\frac{p(\boldsymbol{y})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})}{p(\tilde{\boldsymbol{y}}|\boldsymbol{a})}\right]-J(q(\boldsymbol{y})),$ (179)
with $J(q(\boldsymbol{y}))$ in (III-C).
We prove the following lower bound on the first term in (179).
Lemma 10
Suppose that $p_{\boldsymbol{y}|\boldsymbol{x}}$ is permutation-invariant, i.e. $p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{P}\boldsymbol{y}|\boldsymbol{P}\boldsymbol{x})=p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})$ for any permutation matrix $\boldsymbol{P}$. Then, for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$ we have
$\mathbb{E}\left[\log\frac{p(\boldsymbol{y})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})}{p(\tilde{\boldsymbol{y}}|\boldsymbol{a})}\right]\geq\frac{1}{2}\log\left(1-u_{\mathrm{max}}^{4}\alpha^{2}\right)$
$+\frac{\alpha}{2k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]-\log\left\{1+f(\alpha)\left[(1-k/N)^{-k}-1\right]\right\},$ (180)
with $f(\alpha)\geq 0$ given in (32).
Proof:
See Appendix H. ∎
We are ready to prove Lemma 3. Since the AWGN channel (1) implies the permutation invariance of $p_{\boldsymbol{y}|\boldsymbol{x}}$, we can apply Lemma 10 to the lower bound (179),
$\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}-\hat{\boldsymbol{x}}(\boldsymbol{y})\|_{2}^{2}$
$\geq\frac{(1-u_{\mathrm{max}}^{4}\alpha^{2})e^{\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]}}{\{1+f(\alpha)[(1-k/N)^{-k}-1]\}^{2}}e^{-2J(q(\boldsymbol{y}))}-1$ (181)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$. Using $J(q(\boldsymbol{y}))=D(p(\boldsymbol{y})\|q(\boldsymbol{y}))$ in (III-C) for the Gaussian pdf $q(\boldsymbol{y})$, we arrive at Lemma 3.
Appendix H Proof of Lemma 10
Let $\lambda_{N+1}=\mathbb{E}[y_{N+1}^{2}|\boldsymbol{a}]$ for $y_{N+1}$ in (173), with
$\lambda_{N+1}=1+\frac{\alpha}{k}\boldsymbol{a}^{\mathrm{T}}\mathbb{E}\left[\boldsymbol{x}\boldsymbol{x}^{\mathrm{T}}\right]\boldsymbol{a},$ (182)
where $\boldsymbol{x}$ is uniformly distributed on $\mathcal{X}_{k}^{N}(\mathcal{U})$. In particular, we use $\boldsymbol{a}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{N})$ to have
$\mathbb{E}[\lambda_{N+1}]=1+\frac{\alpha}{k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}].$ (183)
We recall that $\boldsymbol{y}$ given $\boldsymbol{x}$ is represented with the conditional pdf $p_{\boldsymbol{y}|\boldsymbol{x}}$. Using $p(\tilde{\boldsymbol{y}}|\boldsymbol{a})=\mathbb{E}_{\boldsymbol{x}}[p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})]$ and the definition of $y_{N+1}$ in (173) yields
$\mathbb{E}\left[\log\frac{p(\boldsymbol{y})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})}{p(\tilde{\boldsymbol{y}}|\boldsymbol{a})}\right]$
$=\frac{\mathbb{E}[\lambda_{N+1}]-1}{2}+\mathbb{E}[\log p(\boldsymbol{y})]-\mathbb{E}\left[\log\chi(\tilde{\boldsymbol{y}},\boldsymbol{a})\right]$
$=\frac{\alpha}{2k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]+\mathbb{E}[\log p(\boldsymbol{y})]-\mathbb{E}\left[\log\chi(\tilde{\boldsymbol{y}},\boldsymbol{a})\right],$ (184)
with
$\chi(\tilde{\boldsymbol{y}},\boldsymbol{a})=\sum_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{U})}\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime})e^{\sqrt{\alpha/k}\boldsymbol{a}^{\mathrm{T}}\boldsymbol{x}^{\prime}y_{N+1}-\frac{(\alpha/k)(\boldsymbol{a}^{\mathrm{T}}\boldsymbol{x}^{\prime})^{2}}{2}}}{|\mathcal{X}_{k}^{N}(\mathcal{U})|},$ (185)
where the last equality in (184) follows from the expression of $\mathbb{E}[\lambda_{N+1}]$ in (183).
We upper-bound the last term in (184). Substituting the definition of $y_{N+1}$ in (173) and using Jensen’s inequality with respect to $\omega_{N+1}\sim\mathcal{N}(0,1)$ and $\boldsymbol{a}$, we obtain $\mathbb{E}[\log\chi(\tilde{\boldsymbol{y}},\boldsymbol{a})]\leq\mathbb{E}[\log\bar{\chi}(\boldsymbol{y},\boldsymbol{x})]$ with
$\bar{\chi}(\boldsymbol{y},\boldsymbol{x})=\sum_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{U})}\frac{p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime})\mathbb{E}[e^{(\alpha/k)\boldsymbol{a}^{\mathrm{T}}\boldsymbol{x}^{\prime}\boldsymbol{a}^{\mathrm{T}}\boldsymbol{x}}|\boldsymbol{x}]}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}.$ (186)
We repeat the derivation of (99): Using the representation of $\mathcal{X}_{k}^{N}(\mathcal{U})$ in (13) and decomposing $\boldsymbol{x}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S}_{\boldsymbol{x}})$ into $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ in (53) and (54), we have
$\bar{\chi}(\boldsymbol{y},\boldsymbol{x})=\frac{1}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{w=0}^{k}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\bar{\chi}_{w,\boldsymbol{m}}(\boldsymbol{y},\boldsymbol{x}),$ (187)
where $\bar{\chi}_{w,\boldsymbol{m}}(\boldsymbol{y},\boldsymbol{x})$ is given by
$\bar{\chi}_{w,\boldsymbol{m}}=\sum_{\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}}\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime})e_{w,\boldsymbol{m}}(\boldsymbol{x}_{1}^{\prime},\boldsymbol{x}_{2}^{\prime},\boldsymbol{x})$ (188)
for $w\in\{0,\ldots,k\}$, with $\boldsymbol{a}_{1}=\boldsymbol{a}_{\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{a}_{2}=\boldsymbol{a}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{x}_{1}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$, $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}^{\prime}$, $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}^{\prime}$, and
$e_{w,\boldsymbol{m}}=\mathbb{E}\left[\left.\exp\left\{\frac{\alpha}{k}(\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}^{\prime}+\boldsymbol{a}_{2}^{\mathrm{T}}\boldsymbol{x}_{2}^{\prime})\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}\right\}\right|\boldsymbol{x}\right].$ (189)
In particular, for $w=0$ we have $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{0}$ while $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{0}$ holds for $w=k$.
To evaluate the expectation over $\boldsymbol{a}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{N})$, we prove the following proposition:
Proposition 6
For all deterministic $v\in\mathbb{R}$ and $\boldsymbol{u}_{1},\boldsymbol{u}_{2}\in\mathbb{R}^{k}$,
$\mathbb{E}\left[e^{\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{1}(v\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{1}+\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{2})}\right]$
$=\left\{(1-\boldsymbol{u}_{2}^{\mathrm{T}}\boldsymbol{u}_{1})^{2}-(2v+\|\boldsymbol{u}_{2}\|_{2}^{2})\|\boldsymbol{u}_{1}\|_{2}^{2}\right\}^{-1/2}$ (190)
holds if $(1-\boldsymbol{u}_{2}^{\mathrm{T}}\boldsymbol{u}_{1})^{2}>(2v+\|\boldsymbol{u}_{2}\|_{2}^{2})\|\boldsymbol{u}_{1}\|_{2}^{2}$ is satisfied.
Proof:
Let $Z_{i}=\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{i}$ for $i\in\{1,2\}$. The random variables $(Z_{1},Z_{2})$ follow the zero-mean Gaussian distribution with covariance $\mathbb{E}[Z_{i}Z_{j}]=\boldsymbol{u}_{i}^{\mathrm{T}}\boldsymbol{u}_{j}$ for $i,j\in\{1,2\}$. Thus, we can represent $Z_{2}$ as
$Z_{2}=\frac{\boldsymbol{u}_{2}^{\mathrm{T}}\boldsymbol{u}_{1}}{\|\boldsymbol{u}_{1}\|_{2}^{2}}Z_{1}+W,$ (191)
with $W\sim\mathcal{N}(0,\|\boldsymbol{u}_{2}\|_{2}^{2}-(\boldsymbol{u}_{2}^{\mathrm{T}}\boldsymbol{u}_{1})^{2}/\|\boldsymbol{u}_{1}\|_{2}^{2})$ independent of $Z_{1}$. Substituting this representation and evaluating the expectation over $W$, we have
$\mathbb{E}\left[e^{\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{1}(v\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{1}+\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{u}_{2})}\right]$
$=\mathbb{E}\left[\exp\left(\left\{2v+\frac{2\boldsymbol{u}_{2}^{\mathrm{T}}\boldsymbol{u}_{1}}{\|\boldsymbol{u}_{1}\|_{2}^{2}}+\|\boldsymbol{u}_{2}\|_{2}^{2}-\frac{(\boldsymbol{u}_{2}^{\mathrm{T}}\boldsymbol{u}_{1})^{2}}{\|\boldsymbol{u}_{1}\|_{2}^{2}}\right\}\frac{Z_{1}^{2}}{2}\right)\right]$
$=\left\{(1-\boldsymbol{u}_{2}^{\mathrm{T}}\boldsymbol{u}_{1})^{2}-(2v+\|\boldsymbol{u}_{2}\|_{2}^{2})\|\boldsymbol{u}_{1}\|_{2}^{2}\right\}^{-1/2},$ (192)
where the last equality follows from $Z_{1}\sim\mathcal{N}(0,\|\boldsymbol{u}_{1}\|_{2}^{2})$ and the assumption in Proposition 6. ∎
For $w=0$, we have $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{0}$ in $e_{0,\boldsymbol{m}}$ given by (189). we use Proposition 6 with $\boldsymbol{u}_{1}=\sqrt{\alpha/k}\boldsymbol{u}_{\boldsymbol{m}}$, $\boldsymbol{u}_{2}=\sqrt{\alpha/k}\boldsymbol{x}_{1}$, and $v=0$ to have
$e_{0,\boldsymbol{m}}$ $=\left\{[1-(\alpha/k)\boldsymbol{u}_{\boldsymbol{m}}^{\mathrm{T}}\boldsymbol{x}_{1}]^{2}-(\alpha/k)^{2}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}\|\boldsymbol{x}_{1}\|_{2}^{2}\right\}^{-1/2}$
$\leq\left(1-2u_{\mathrm{max}}^{2}\alpha\right)^{-1/2}$ (193)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$, where the last inequality follows from the upper bounds $\boldsymbol{u}_{\boldsymbol{m}}^{\mathrm{T}}\boldsymbol{x}_{1}\leq u_{\mathrm{max}}^{2}k$, $\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}\leq u_{\mathrm{max}}^{2}k$, and $\|\boldsymbol{x}_{1}\|_{2}^{2}\leq u_{\mathrm{max}}^{2}k$ for all $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$.
For $w=k$, we have $\boldsymbol{x}_{1}^{\prime}=\boldsymbol{0}$ in $e_{k,\boldsymbol{m}}$. Evaluating the expectation over $\boldsymbol{a}_{2}$ and subsequently over $\boldsymbol{a}_{1}$ yields
$e_{k,\boldsymbol{m}}=\mathbb{E}\left[\left.e^{\frac{\alpha^{2}}{2k^{2}}\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}(\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1})^{2}}\right|\boldsymbol{x}\right]$
$=\left(1-\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}\|\boldsymbol{x}_{1}\|_{2}^{2}\alpha^{2}/k^{2}\right)^{-1/2}\leq\left(1-u_{\mathrm{max}}^{4}\alpha^{2}\right)^{-1/2}$ (194)
for all $\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{k,\boldsymbol{m}}^{N-k,2}$ and $\alpha\in(0,u_{\mathrm{max}}^{-2})$.
For $w\in\{1,\ldots,k-1\}$ we evaluate the expectation in (189) over $\boldsymbol{a}_{2}$ as
$e_{w,\boldsymbol{m}}=\mathbb{E}\left[\left.\exp\left(\frac{\alpha\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}^{\prime}\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1}}{k}+\frac{\alpha^{2}\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}(\boldsymbol{a}_{1}^{\mathrm{T}}\boldsymbol{x}_{1})^{2}}{2k^{2}}\right)\right|\boldsymbol{x}\right].$ (195)
Using Proposition 6 with $\boldsymbol{u}_{1}=(\alpha/k)\boldsymbol{x}_{1}$, $\boldsymbol{u}_{2}=\boldsymbol{x}_{1}^{\prime}$, and $v=\|\boldsymbol{u}_{\boldsymbol{m}_{2}}\|_{2}^{2}/2$ yields
$e_{w,\boldsymbol{m}}$ $=\left\{[1-(\alpha/k)\boldsymbol{x}_{1}^{\prime\mathrm{T}}\boldsymbol{x}_{1}]^{2}-(\alpha^{2}/k^{2})\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}\|\boldsymbol{x}_{1}\|_{2}^{2}\right\}^{-1/2}$
$\leq\left(1-2u_{\mathrm{max}}^{2}\alpha\right)^{-1/2}$ (196)
for all $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ and $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$, where the last inequality follows from the upper bounds $\boldsymbol{x}_{1}^{\prime\mathrm{T}}\boldsymbol{x}_{1}\leq u_{\mathrm{max}}^{2}k$, $\|\boldsymbol{u}_{\boldsymbol{m}}\|_{2}^{2}\leq u_{\mathrm{max}}^{2}k$, and $\|\boldsymbol{x}_{1}\|_{2}^{2}\leq u_{\mathrm{max}}^{2}k$ for all $\boldsymbol{x}\in\mathcal{X}_{k}^{N}(\mathcal{U})$ and $\boldsymbol{x}_{1}^{\prime}\in\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$. Applying these results to (187), we have
$\bar{\chi}(\boldsymbol{y},\boldsymbol{x})\leq$ $(1-2u_{\mathrm{max}}^{2}\alpha)^{-1/2}\left\{p(\boldsymbol{y})-\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})\right\}$
$+(1-u_{\mathrm{max}}^{4}\alpha^{2})^{-1/2}\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y}),$ (197)
with
$\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})=\frac{1}{|\mathcal{X}_{k}^{N}(\mathcal{U})|}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\sum_{\boldsymbol{x}_{2}^{\prime}\in\mathcal{T}_{k,\boldsymbol{m}}^{N-k,2}}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime}),$ (198)
with $\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}^{\prime}=\boldsymbol{0}$ and $\boldsymbol{x}_{2}^{\prime}=\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}^{\prime}$. The quantity $\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})$ depends on $\mathcal{S}_{\boldsymbol{x}}$ only through the partition of $\boldsymbol{x}^{\prime}$.
We evaluate the quantity (184). Applying (197) and the inequality $\mathbb{E}[\log\chi(\tilde{\boldsymbol{y}},\boldsymbol{a})]\leq\mathbb{E}[\log\bar{\chi}(\boldsymbol{y},\boldsymbol{x})]$ to (184), we obtain
$\mathbb{E}\left[\log\frac{p(\boldsymbol{y})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})}{p(\tilde{\boldsymbol{y}}|\boldsymbol{a})}\right]\geq\frac{1}{2}\log\left(1-u_{\mathrm{max}}^{4}\alpha^{2}\right)$
$-\mathbb{E}\left[\log\left\{1+f(\alpha)\left[1-\frac{\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})}{p(\boldsymbol{y})}\right]\right\}\right]+\frac{\alpha}{2k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]$ (199)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$, with $f(\alpha)\geq 0$ given in (32). Using Jensen’s inequality for the second term in (199) yields
$\mathbb{E}\left[\log\left\{1+f(\alpha)\left[1-\frac{\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})}{p(\boldsymbol{y})}\right]\right\}\right]$
$\leq\log\left(1+f(\alpha)\left\{1-\mathbb{E}\left[\frac{\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})}{p(\boldsymbol{y})}\right]\right\}\right).$ (200)
To evaluate the expectation in (200), we utilize the permutation-invariance of $p(\boldsymbol{y}|\boldsymbol{S}_{\boldsymbol{x}})$—written as $p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{y}|\boldsymbol{S}_{\boldsymbol{x}})$ for clarification, given by
$p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{y}|\mathcal{S})=\frac{1}{|\mathcal{M}|^{k}}\sum_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{S},\mathcal{U})}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime}),$ (201)
with $\mathcal{X}_{k}^{N}(\mathcal{S},\mathcal{U})=\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{U}):\mathcal{S}_{\boldsymbol{x}^{\prime}}=\mathcal{S}\}$ for $\mathcal{S}\in\mathfrak{S}_{k}^{N}$. By definition, we have $|\mathcal{X}_{k}^{N}(\mathcal{S},\mathcal{U})|=|\mathcal{M}|^{k}$.
The permutation invariance of $p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}$ is defined via any $N\times N$ permutation matrix $\boldsymbol{P}$, which is a bijection from $\mathcal{X}_{k}^{N}(\{1\})$ onto $\mathcal{X}_{k}^{N}(\{1\})$. We define the one-to-one mapping $\pi_{\boldsymbol{P}}:\mathfrak{S}_{k}^{N}\to\mathfrak{S}_{k}^{N}$ induced from $\boldsymbol{P}$ as
$\boldsymbol{s}^{\prime}=\boldsymbol{P}\boldsymbol{s}\iff\mathcal{S}_{\boldsymbol{s}^{\prime}}=\pi_{\boldsymbol{P}}(\mathcal{S}_{\boldsymbol{s}})$ (202)
for all $\boldsymbol{s},\boldsymbol{s}^{\prime}\in\mathcal{X}_{k}^{N}(\{1\})$. Then, for all $\boldsymbol{y}\in\mathbb{R}^{N}$ and $\mathcal{S}\in\mathfrak{S}_{k}^{N}$ we have the following permutation invariance:
$p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{P}\boldsymbol{y}|\pi_{\boldsymbol{P}}(\mathcal{S}))=\frac{1}{|\mathcal{M}|^{k}}\sum_{\boldsymbol{x}^{\prime\prime}\in\mathcal{X}_{k}^{N}(\pi_{\boldsymbol{P}}(\mathcal{S}),\mathcal{U})}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{P}\boldsymbol{y}|\boldsymbol{x}^{\prime\prime})$
$=\frac{1}{|\mathcal{M}|^{k}}\sum_{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{S},\mathcal{U})}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{P}\boldsymbol{y}|\boldsymbol{P}\boldsymbol{x}^{\prime})=p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{y}|\mathcal{S}).$ (203)
In the derivation of the second equality, we have used the change of variables $\boldsymbol{x}^{\prime}=\boldsymbol{P}^{\mathrm{T}}\boldsymbol{x}^{\prime\prime}$ and $\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{S},\mathcal{U})$, of which the latter is obtained from the definition of $\pi_{\boldsymbol{P}}$. The last equality follows from the permutation-invariance assumption of $p_{\boldsymbol{y}|\boldsymbol{x}}$ in Lemma 10.
Similarly, we confirm the permutation invariance of $p(\boldsymbol{y})$—written as $p_{\boldsymbol{y}}(\boldsymbol{y})$. Since $\mathcal{S}_{\boldsymbol{x}}$ is uniformly distributed on $\mathfrak{S}_{k}^{N}$, we have the representation $p_{\boldsymbol{y}}(\boldsymbol{y})=|\mathfrak{S}_{k}^{N}|^{-1}\sum_{\mathcal{S}\in\mathfrak{S}_{k}^{N}}p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{y}|\mathcal{S})$. Repeating the proof (203) with the permutation invariance of $p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}$, we find the permutation invariance of $p_{\boldsymbol{y}}(\boldsymbol{y})$.
We are ready to evaluate the expectation in (200). Let $\mathcal{S}^{\prime}$ denote the support of $\boldsymbol{x}_{2}^{\prime}$ in (198). By definition, $\mathcal{S}^{\prime}$ is in $\{\mathcal{S}^{\prime}\in\mathfrak{S}_{k}^{N}:\mathcal{S}\cap\mathcal{S}^{\prime}=\emptyset\}$ for given $\mathcal{S}_{\boldsymbol{x}}=\mathcal{S}$. Since $\mathcal{S}_{\boldsymbol{x}}$ is uniformly distributed on $\mathfrak{S}_{k}^{N}$, we use the definition of $\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})$ in (198) to have
$\mathbb{E}\left[\frac{\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})}{p_{\boldsymbol{y}}(\boldsymbol{y})}\right]=\frac{1}{|\mathfrak{S}_{k}^{N}|}\sum_{\mathcal{S}\in\mathfrak{S}_{k}^{N}}\int\frac{\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})}{p_{\boldsymbol{y}}(\boldsymbol{y})}p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{y}|\mathcal{S})d\boldsymbol{y}$
$=\sum_{\mathcal{S}\in\mathfrak{S}_{k}^{N}}\sum_{\mathcal{S}^{\prime}\in\mathfrak{S}_{k}^{N}:\mathcal{S}\cap\mathcal{S}^{\prime}=\emptyset}\frac{f(\mathcal{S},\mathcal{S}^{\prime})}{|\mathfrak{S}_{k}^{N}||\mathcal{X}_{k}^{N}(\mathcal{U})|},$ (204)
where $f(\mathcal{S},\mathcal{S}^{\prime})$ is given by
$f(\mathcal{S},\mathcal{S}^{\prime})=\int\frac{p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{y}|\mathcal{S})}{p_{\boldsymbol{y}}(\boldsymbol{y})}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime})d\boldsymbol{y},$ (205)
with $\boldsymbol{x}_{\mathcal{S}^{\prime}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}$ and $\boldsymbol{x}_{\mathcal{N}\setminus\mathcal{S}^{\prime}}^{\prime}=\boldsymbol{0}$.
We confirm that the function (205) does not depend on $\mathcal{S}^{\prime}$. Let $\bar{\mathcal{S}}^{\prime}\in\mathfrak{S}_{k}^{N}$ with $\mathcal{S}\cap\bar{\mathcal{S}}^{\prime}=\emptyset$ denote a reference support. We write the set of the corresponding reference vectors as $\bar{\mathcal{X}}^{\prime}(\bar{\mathcal{S}}^{\prime})=\{\bar{\boldsymbol{x}}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{U}):\bar{\boldsymbol{x}}_{\bar{\mathcal{S}}^{\prime}}^{\prime}\in\mathcal{M}^{k}\}$. For any $\mathcal{S}^{\prime}\in\mathfrak{S}_{k}^{N}$ with $\mathcal{S}\cap\mathcal{S}^{\prime}=\emptyset$, there is some permutation matrix $\boldsymbol{P}$ such that $\boldsymbol{P}$ is a bijection from $\bar{\mathcal{X}}^{\prime}(\mathcal{S}^{\prime})$ onto $\bar{\mathcal{X}}^{\prime}(\bar{\mathcal{S}}^{\prime})$. The corresponding mapping $\pi_{\boldsymbol{P}}$ satisfies $\pi_{\boldsymbol{P}}(\mathcal{S}^{\prime})=\bar{\mathcal{S}}^{\prime}$. When $\mathcal{S}$ has no intersection with $\mathcal{S}^{\prime}$, we find the identity $\pi_{\boldsymbol{P}}(\mathcal{S})=\mathcal{S}$. Using the permutation invariance of the pdfs $p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}$, $p_{\boldsymbol{y}}$, and $p_{\boldsymbol{y}|\boldsymbol{x}}$, as well as $\pi_{\boldsymbol{P}}(\mathcal{S})=\mathcal{S}$, we evaluate (205) as
$f(\mathcal{S},\mathcal{S}^{\prime})=\int\frac{p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{P}\boldsymbol{y}|\mathcal{S})}{p_{\boldsymbol{y}}(\boldsymbol{P}\boldsymbol{y})}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{P}\boldsymbol{y}|\boldsymbol{P}\boldsymbol{x}^{\prime})d\boldsymbol{y}$
$=\int\frac{p_{\boldsymbol{y}|\mathcal{S}_{\boldsymbol{x}}}(\boldsymbol{y}|\mathcal{S})}{p_{\boldsymbol{y}}(\boldsymbol{y})}\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\bar{\boldsymbol{x}}^{\prime})d\boldsymbol{y}=f(S,\bar{S}^{\prime}),$ (206)
with $\bar{\boldsymbol{x}}_{\bar{\mathcal{S}^{\prime}}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}$ and $\bar{\boldsymbol{x}}_{\mathcal{N}\setminus\bar{\mathcal{S}^{\prime}}}^{\prime}=\boldsymbol{0}$, where the second equality follows from the change of variables in the integral and the definition of the permutation matrix $\boldsymbol{P}$.
We evaluate the quantity (204). Taking the average of $f(\mathcal{S},\bar{\mathcal{S}}^{\prime})$ in (205) over $\mathcal{S}\in\mathfrak{S}_{k}^{N}$, we use the definition of $p_{\boldsymbol{y}}(\boldsymbol{y})$ to have
$\frac{1}{|\mathfrak{S}_{k}^{N}|}\sum_{\mathcal{S}\in\mathfrak{S}_{k}^{N}}f(\mathcal{S},\bar{\mathcal{S}}^{\prime})=\sum_{\boldsymbol{m}\in\mathcal{M}^{k}}\int p_{\boldsymbol{y}|\boldsymbol{x}}(\boldsymbol{y}|\boldsymbol{x}^{\prime})d\boldsymbol{y}=|\mathcal{M}|^{k}.$ (207)
From this identity and (206), we evaluate (204) as
$\mathbb{E}\left[\frac{\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})}{p_{\boldsymbol{y}}(\boldsymbol{y})}\right]$ $=\binom{N-k}{k}\sum_{\mathcal{S}\in\mathfrak{S}_{k}^{N}}\frac{f(\mathcal{S},\bar{\mathcal{S}}^{\prime})}{|\mathfrak{S}_{k}^{N}||\mathcal{X}_{k}^{N}(\mathcal{U})|}$
$=\binom{N-k}{k}\binom{N}{k}^{-1},$ (208)
because of $|\mathcal{X}_{k}^{N}(\mathcal{U})|=|\mathcal{M}|^{k}\binom{N}{k}$. Using $\sum_{w=0}^{k}\binom{k}{w}\binom{N-k}{w}=\binom{N}{k}$, we arrive at
$1-\mathbb{E}\left[\frac{\Delta_{\mathcal{S}_{\boldsymbol{x}}}p(\boldsymbol{y})}{p_{\boldsymbol{y}}(\boldsymbol{y})}\right]=\binom{N}{k}^{-1}\sum_{w=0}^{k-1}\binom{k}{w}\binom{N-k}{w}$
$\leq\{1+(N/k-1)^{-1}\}^{k}-1=(1-k/N)^{-k}-1,$ (209)
where the inequality follows from the upper bound (105) and the binomial theorem.
We are ready to prove Lemma 10. Applying (200) and (209) to (199), we obtain
$\mathbb{E}\left[\log\frac{p(\boldsymbol{y})p(y_{N+1}|\boldsymbol{x},\boldsymbol{a})}{p(\tilde{\boldsymbol{y}}|\boldsymbol{a})}\right]\geq\frac{1}{2}\log\left(1-u_{\mathrm{max}}^{4}\alpha^{2}\right)$
$+\frac{\alpha}{2k}\mathbb{E}[\|\boldsymbol{x}\|_{2}^{2}]-\log\left\{1+f(\alpha)\left[(1-k/N)^{-k}-1\right]\right\}$ (210)
for all $\alpha\in(0,(2u_{\mathrm{max}}^{2})^{-1})$.
Acknowledgment
The author thanks the anonymous reviewers for their suggestions that have improved the quality of the manuscript greatly.
TABLE I: List of notation
Notation Definition
$\gamma\in[0,1)$ $\log k/\log N\to\gamma$
$\mathcal{N}$ $\{1,\ldots,N\}$
$\mathcal{M}$ $\{1,\ldots,M\}$
$\mathfrak{S}_{k}^{N}$ $\{\mathcal{S}\subset\mathcal{N}:|\mathcal{S}|=k\}$
$\mathcal{S}_{\boldsymbol{x}}$ $\{n\in\mathcal{N}:x_{n}\neq 0\}$
$\mathcal{U}$ $\{u_{m}\in\mathbb{R}\setminus\{0\}:m\in\mathcal{M}\}$
$\mathcal{X}_{k}^{N}(\mathcal{U})$ $\{\boldsymbol{x}\in\mathbb{R}^{N}:|\mathcal{S}_{\boldsymbol{x}}|=k,x_{n}\in\mathcal{U}~\hbox{for all $n\notin\mathcal{S}_{\boldsymbol{x}}$}\}$
$\mathcal{X}_{k}^{N}(\mathcal{S},\mathcal{U})$ $\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k}^{N}(\mathcal{U}):\mathcal{S}_{\boldsymbol{x}^{\prime}}=\mathcal{S}\}$ for $\mathcal{S}\in\mathfrak{S}_{k}^{N}$
$u_{\mathrm{min}}$ $\min_{m\in\mathcal{M}}|u_{m}|$
$u_{\mathrm{max}}$ $\max_{m\in\mathcal{M}}|u_{m}|$
$d_{\mathrm{max}}$ $\max_{m,m^{\prime}\in\mathcal{M}:m\neq m^{\prime}}|u_{m}-u_{m^{\prime}}|$
$\sigma_{N/k}^{2}$ $\sigma^{2}/\log(N/k)$
$N_{\boldsymbol{x},\hat{\boldsymbol{x}}_{\mathrm{ML}}}$ $|\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}|-|\{n\in\mathcal{S}_{\boldsymbol{x}}\cap\mathcal{S}_{\hat{\boldsymbol{x}}_{\mathrm{ML}}}:x_{n}=\hat{x}_{\mathrm{ML},n}\}|$
$\boldsymbol{u}_{\boldsymbol{m}}\in\mathcal{U}^{k}$ $u_{m_{1}},\ldots,u_{m_{k}}]^{\mathrm{T}}$
$\mathcal{T}_{w,\boldsymbol{m}}^{N}(\mathcal{S})$ $\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{|\mathcal{S}|}^{N}(\mathcal{U}):|\mathcal{S}_{\boldsymbol{x}^{\prime}}\setminus\mathcal{S}|=w,\;\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}^{\prime}}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}}\}$
$\mathcal{T}_{k-w,\boldsymbol{m}_{1}}^{k,1}$ $\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{k-w}^{k}(\mathcal{U}):\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}^{\prime}}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}_{1}}\}$ for $\boldsymbol{m}_{1}\in\mathcal{M}^{k-w}$
$\mathcal{T}_{w,\boldsymbol{m}_{2}}^{N-k,2}$ $\{\boldsymbol{x}^{\prime}\in\mathcal{X}_{w}^{N-k}(\mathcal{U}):\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}^{\prime}}}^{\prime}=\boldsymbol{u}_{\boldsymbol{m}_{2}}\}$ for $\boldsymbol{m}_{2}\in\mathcal{M}^{w}$
$\boldsymbol{x}_{1}$ $\boldsymbol{x}_{\mathcal{S}_{\boldsymbol{x}}}$
$\boldsymbol{\omega}_{1}$ $\boldsymbol{\omega}_{\mathcal{S}_{\boldsymbol{x}}}$
$\boldsymbol{\omega}_{2}$ $\boldsymbol{\omega}_{\mathcal{N}\setminus\mathcal{S}_{\boldsymbol{x}}}$
$\Omega_{i}^{w}$ $(\boldsymbol{x}_{i}^{w})^{\mathrm{T}}\boldsymbol{\omega}_{1}/\sigma_{N/k}^{2}$ for $\boldsymbol{x}_{i}^{w}\in\mathcal{X}_{k-w}^{k}(\mathcal{U})$
$d_{i}^{w}$ $u_{\mathrm{max}}^{-1}\|\boldsymbol{x}_{i}^{w}\|_{1}d_{*}$ with $d_{*}$ in (110)
$\mathcal{E}_{i}^{w}$ $\{\Omega_{i}^{w}:|\Omega_{i}^{w}|\leq d_{i}^{w}\}$
$\mathcal{E}_{\mathrm{b}}$ $\cap_{i=1}^{k}\mathcal{E}_{i}^{k-1}$